### v1

# Generalized quantum Zernike Hamiltonians: Polynomial Higgs-type algebras and algebraic derivation of the spectrum

Abstract
We consider the quantum analog of the generalized Zernike systems given by the Hamiltonian:
^$\hat{\mathcal{H}}_{N}=\hat{p}_{1}^{2}+\hat{p}_{2}^{2}+\sum_{k=1}^{N}\gamma_{k}(\hat{q}_{1}\hat{p}_{1}+\hat{q}_{2}\hat{p}_{2})^{k},$
with canonical operators $\hat{q}_{i},\,\hat{p}_{i}$ and arbitrary coefficients $\gamma_{k}$. This two-dimensional quantum model, besides the conservation of the angular momentum, exhibits higher-order integrals of motion within the enveloping algebra of the Heisenberg algebra $\mathfrak{h}_{2}$. By constructing suitable combinations of these integrals, we uncover a polynomial Higgs-type symmetry algebra that, through an appropriate change of basis, gives rise to a deformed oscillator algebra. The associated structure function $\Phi$ is shown to factorize into two commuting components $\Phi=\Phi_{1}\Phi_{2}$. This framework enables an algebraic determination of the possible energy spectra of the model for the cases $N=2,3,4$, the case $N=1$ being canonically equivalent to the harmonic oscillator. Based on these findings, we propose two conjectures which generalize the results for all $N\geq 2$ and any value of the coefficients $\gamma_{k}$, that they are explicitly proven for $N=5$. In addition, all of these results can be interpreted as superintegrable perturbations of the original quantum Zernike system corresponding to $N=2$ which are also analyzed and applied to the isotropic oscillator on the sphere, hyperbolic and Euclidean spaces.
To the memory of our colleague and friend Prof. George Samvelovich Pogosyan
Keywords: Zernike system; superintegrability; deformed oscillator algebras; polynomial symmetry algebras; eigenvalue spectra; momentum-dependent potentials; integrable perturbations; curved oscillators
PACS: 02.30.Ik, 45.20.Jj, 02.20.Sv, 02.40.Ky
MSC: 37J35, 22E60, 46N50
1 Introduction
The so-called Zernike system, emerged in the context of wavefront descriptions in optics [2], has turned out to be an extremely interesting model, with many physically and mathematically relevant applications. Besides being the first case in a hierarchy of (classically) superintegrable systems [3], its quantum version determines a new class of orthogonal polynomials, the Zernike polynomials, of great interest in the treatment of optical aberrations. In addition, it has been shown that the quantum symmetries of the system give rise to a cubic Higgs algebra $\mathfrak{sl}^{(3)}(2,\mathbb{R})$. The latter fact, in combination with recent work on polynomial algebras in connection with the superintegrability property (see e.g. [4] and references therein), suggest to look for a quantization of the generalized classical Zernike Hamiltonians considered in [3] in such a way that the corresponding quantum symmetry algebra is a higher-order polynomial Higgs-type algebra.
The two-dimensional Zernike system is defined by means of the following quantum Hamiltonian ^$\hat{\mathcal{H}}_{\rm Zk}$ with eigenvalue equation (with normalized units $\hbar=1$)
^$\hat{\mathcal{H}}_{\rm Zk}\psi(\mathbf{q})=\left(-\nabla^{2}-\beta\,{\mathbf{q}}\boldsymbol{\cdot}\nabla-\alpha\bigl{(}{\mathbf{q}}\boldsymbol{\cdot}\nabla\bigr{)}^{2}\right)\psi(\mathbf{q})=E\,\psi(\mathbf{q}),$ (1.1)
where we have denoted
${\mathbf{q}}=(q_{1},q_{2}),\qquad\nabla=\left(\frac{\partial}{\partial q_{1}},\frac{\partial}{\partial q_{2}}\right),\qquad{\mathbf{q}}\boldsymbol{\cdot}\nabla=q_{1}\frac{\partial}{\partial q_{1}}+q_{2}\frac{\partial}{\partial q_{2}},$
such that $\alpha$ and $\beta$ are two real parameters. It turns out that the Zernike system exhibits interesting structural features of great relevance in quantum optics, which have motivated its detailed study in recent years (see [5, 6, 7, 8, 9, 10] and references therein). Among these important properties, the Zernike system is superintegrable, that is, it admits two algebraically independent quantum observables that commute with the Hamiltonian ^$\hat{\mathcal{H}}_{\rm Zk}$. Therefore, the corresponding spectrum presents a maximal degeneracy. We also recall that, as shown in [2], among the possible values for the real parameters $\alpha$ and $\beta$, the case with $\alpha=-1$ and $\beta=-2$ is the only one that provides a self-adjoint operator ^$\hat{\mathcal{H}}_{\rm Zk}$ under the inner product over the unit disk with $|{\mathbf{q}}|\leq 1$, which has been fully studied by Pogosyan et al in [5], leading to the Zernike polynomials (see e.g. [11, 12] and references therein).
On the other hand, the classical Hamiltonian counterpart of ^$\hat{\mathcal{H}}_{\rm Zk}$ (1.1), also considered by Pogosyan et al in [13], adopts the following form
$\mathcal{H}_{\rm Zk}=p_{1}^{2}+p_{2}^{2}-{\rm i}\beta(q_{1}p_{1}+q_{2}p_{2})+\alpha(q_{1}p_{1}+q_{2}p_{2})^{2},$ (1.2)
where $q_{i}$ and $p_{i}$ are generic canonical variables fulfilling the Poisson bracket $\{q_{i},p_{j}\}=\delta_{ij}$. Obviously, the Hamiltonian $\mathcal{H}_{\rm Zk}$ determines a classical superintegrable system, meaning that it admits two constants of the motion, in this case quadratic in the momenta, which added to the Hamiltonian $\mathcal{H}_{\rm Zk}$ form a set of three functionally independent functions. It follows that all bounded trajectories are periodic, and for $\mathcal{H}_{\rm Zk}$ these are given by ellipses [13]. A first superintegrable generalization of the classical Zernike Hamiltonian (1.2) was considered and analyzed in [3] in the form
$\mathcal{H}_{N}=p_{1}^{2}+p_{2}^{2}+\sum_{k=1}^{N}\gamma_{k}(q_{1}p_{1}+q_{2}p_{2})^{k},$ (1.3)
where $\gamma_{k}$ are arbitrary coefficients that can be either pure imaginary or real numbers. Two constants of the motion for the Hamiltonian $\mathcal{H}_{N}$ are the usual angular momentum
$\mathcal{C}=q_{1}p_{2}-q_{2}p_{1}$ (1.4)
and a certain function $\mathcal{I}_{N}$, which is of $N^{{\rm th}}$-order in the momenta and has been obtained explicitly in [3] for arbitrary index $N$. The classical Zernike system $\mathcal{H}_{\rm Zk}$ is recovered from $\mathcal{H}_{N}$ (1.3) for the quadratic case with $N=2$ by setting
$\gamma_{1}=-\rm i\beta,\qquad\gamma_{2}=\alpha,\qquad\alpha,\beta\in\mathbb{R}.$ (1.5)
In this context, $\mathcal{H}_{N}$ can be interpreted as a superintegrable perturbation of $\mathcal{H}_{\rm Zk}$ for $N\geq 3$, and such that the bounded trajectories are no longer ellipses, but some ‘deformation’ of them that increases with the values of $\gamma_{k}$ ($k\geq 3)$ as shown in [3]. More recently, an additional generalization of the Hamiltonian (1.3) has been proposed in [14] in terms of generic analytical functions $F\left(q_{1}p_{1}+q_{2}p_{2}\right)$, showing that the superintegrability property actually arises from the particular structure of the Hamiltonian.
The aim of this work is twofold. In the first place, we propose a generalized quantum Zernike Hamiltonian ^$\hat{\mathcal{H}}_{N}$ by imposing that the following requirements are satisfied:
(i) The superintegrability property is preserved for any $N$.
(ii) For $N=2$, the quantum Zernike system ^$\hat{\mathcal{H}}_{\rm Zk}$ in (1.1) is recovered.
(iii) The classical counterpart of ^$\hat{\mathcal{H}}_{N}$ is given by $\mathcal{H}_{N}$ in (1.3).
As second objective, we derive the corresponding spectrum of ^$\hat{\mathcal{H}}_{\rm Zk}$ by means of an algebraic procedure that involves the quantum symmetries of ^$\hat{\mathcal{H}}_{N}$, i.e., algebraically independent quantum observables commuting with ^$\hat{\mathcal{H}}_{N}$, which are shown to be of $N^{{\rm th}}$-order in the momenta.
Furthermore, we recall that the classical Zernike Hamiltonian $\mathcal{H}_{\rm Zk}$ (1.2) has been alternatively interpreted in [3] as the two-dimensional isotropic oscillator on the sphere, hyperbolic and Euclidean spaces (the former being just the Higgs oscillator), such that the parameter $\beta$ is related to the frequency, while $\alpha$ determines the constant curvature of the underlying space. As a direct consequence, for $N\geq 3$ our results provide superintegrable perturbations of such (curved) quantum oscillators that we also study in detail.
The structure of the paper is as follows: In section 2 we present the algebraic procedure used to obtain the quantum symmetries and spectrum for the generalized quantum Zernike Hamiltonian ^$\hat{\mathcal{H}}_{N}$ in (1.3). In section 3, our approach is first applied to the proper (quadratic) quantum Zernike Hamiltonian ^$\hat{\mathcal{H}}_{\rm Zk}$ (1.1), recovering known results. As a byproduct, the associated quantum curved oscillators are also deduced and studied. Sections 4 and 5 are devoted to the cubic $N=3$ and quartic $N=4$ systems, which are solved explicitly. In addition, we stress that the cubic case is shown to introduce a superintegrable perturbation of the quantum isotropic oscillator on the above three Riemannian spaces that are fully analyzed in section 4.2. Besides the two curved oscillators, this application also yields new quantum Euclidean oscillators which are named as either cubic spherical or hyperbolic perturbations.
Due to ordering problems in the quantization, the computational difficulties increase exponentially for higher values of $N$, so that general expression for arbitrary $N$ remains as an open problem; in fact, even in the classical framework [3], the general construction is quite involved. Nevertheless, we present in section 6 two conjectures for the spectrum of ^$\hat{\mathcal{H}}_{N}$ (1.3) with arbitrary $N$ along with the solution for $N=5$. The paper concludes with some remarks and some open problems whose detailed analysis will be considered in the future.
2 Generalized quantum Zernike Hamiltonians
Let us consider the usual quantum position ${\hat{\mathbf{q}}}=({\hat{q}}_{1},{\hat{q}}_{2})$ and momenta ${\hat{\mathbf{p}}}=({\hat{p}}_{1},{\hat{p}}_{2})$ operators, with canonical Lie brackets and differential realization given by
$[\hat{q}_{i},\hat{q}_{j}]=[\hat{p}_{i},\hat{p}_{j}]=0,\qquad[\hat{q}_{i},\hat{p}_{j}]={\rm i}\,{\rm Id}\,\delta_{ij},\qquad\hat{q}_{i}\psi(\mathbf{q})=q_{i}\psi(\mathbf{q}),\qquad\hat{p}_{i}\psi(\mathbf{q})=-{\rm i}\,\frac{\partial\psi(\mathbf{q})}{\partial q_{i}},$ (2.1)
where ${\rm Id}$ denotes the identity operator and, for clarity in the exposition, we normalize units by setting $\hbar=1$. Under this realization, the quantum Zernike Hamiltonian (1.1) is written as
^$\begin{split}\hat{\mathcal{H}}_{\rm Zk}&={\hat{p}}_{1}^{2}+{\hat{p}}_{2}^{2}-{\rm i}\beta({\hat{q}}_{1}{\hat{p}}_{1}+{\hat{q}}_{2}{\hat{p}}_{2})+\alpha({\hat{q}}_{1}{\hat{p}}_{1}+{\hat{q}}_{2}{\hat{p}}_{2})^{2}\\[2.0pt] &={\hat{\mathbf{p}}}^{2}-{\rm i}\beta({\hat{\mathbf{q}}}\boldsymbol{\cdot}{\hat{\mathbf{p}}})+\alpha({\hat{\mathbf{q}}}\boldsymbol{\cdot}{\hat{\mathbf{p}}})^{2}.\end{split}$ (2.2)
Taking into account the latter expression and the classical superintegrable Hamiltonian $\mathcal{H}_{N}$ (1.3) we shall consider, as a suitable quantum analogue, the following quantum Hamiltonian [3]
^$\hat{\mathcal{H}}_{N}={\hat{\mathbf{p}}}^{2}+\sum_{k=1}^{N}\gamma_{k}({\hat{\mathbf{q}}}\boldsymbol{\cdot}{\hat{\mathbf{p}}})^{k},$ (2.3)
where $\gamma_{k}$ are arbitrary (real or pure imaginary) coefficients. This leads to the Schrödinger equation given by
^$\begin{split}\hat{\mathcal{H}}_{N}\psi(\mathbf{q})&=\left(-\nabla^{2}+\sum_{k=1}^{N}\gamma_{k}(-{\rm i})^{k}\bigl{(}{\mathbf{q}}\boldsymbol{\cdot}\nabla\bigr{)}^{k}\right)\psi(\mathbf{q})=E\,\psi(\mathbf{q})\\[4.0pt] &=\left(-\left(\frac{\partial^{2}}{\partial q_{1}^{2}}+\frac{\partial^{2}}{\partial q_{2}^{2}}\right)+\sum_{k=1}^{N}\gamma_{k}(-{\rm i})^{k}\left(q_{1}\frac{\partial}{\partial q_{1}}+q_{2}\frac{\partial}{\partial q_{2}~{}}\right)^{\!k}\right)\psi(\mathbf{q})=E\,\psi(\mathbf{q}),\end{split}$ (2.4)
which generalizes to arbitrary values of $N\geq 3$ the Hamiltonian ^$\hat{\mathcal{H}}_{\rm Zk}$ (1.1), that is recovered with the choices (1.5). It is important to observe that, due to the term ordering problems arising in the quantization of $\mathcal{H}_{N}$ (1.3) as well as for the corresponding constants of the motion, there are many possibilities for the quantum construction. Nevertheless, we will show here that the Hamiltonian ^$\hat{\mathcal{H}}_{N}$ in (2.3) determines a quantum superintegrable system by explicitly finding its quantum symmetries and then computing its spectrum for arbitrary coefficients $\gamma_{k}$. As expected, for increasing values of $N$, some of the computations become extremely cumbersome, requiring the use of symbolic computer packages, such as the Wolfram Mathematica® software system.
The first step in our construction is to obtain the quantum symmetries of ^$\hat{\mathcal{H}}_{N}$ ensuring the superintegrability property. One of these symmetries clearly corresponds to the quantum angular momentum operator
$\hat{\mathcal{C}}={\hat{q}}_{1}{\hat{p}}_{2}-{\hat{q}}_{2}{\hat{p}}_{1},$ (2.5)
as it commutes with ${\hat{\mathbf{q}}}\boldsymbol{\cdot}{\hat{\mathbf{p}}}$. This shows that ^$\hat{\mathcal{H}}_{N}$ is a quantum integrable Hamiltonian. Due to the structure of the Hamiltonian, we expect the additional quantum symmetry for ^$\hat{\mathcal{H}}_{N}$ to be of higher-order in the momenta. Following this assumption, we consider the enveloping algebra of the Heisenberg algebra in two dimensions spanned by ${\hat{\mathbf{q}}}$ and ${\hat{\mathbf{p}}}$ (2.1), to which the identity operator ${\rm Id}$ is added as fifth generator, and introduce the following grading
$|{\hat{q}}_{1}|=|{\hat{q}}_{2}|=1,\qquad|{\hat{p}}_{1}|=|{\hat{p}}_{2}|=-1.$
Next, we look for monomials $\hat{\mathcal{P}}_{\ell}^{(s)}({\hat{\mathbf{q}}},{\hat{\mathbf{p}}})$ of degree $\ell\leq 2N$ such that the total grading of the monomial is equal to 0. The grading is assigned via the prescription
$|{\hat{q}}_{1}^{a}\,{\hat{q}}_{2}^{b}\,{\hat{p}}_{1}^{c}\,{\hat{p}}_{2}^{d}|=a+b-c-d.$
In this way, we construct the following operators of $N^{{\rm th}}$-order in the momenta
${\hat{I}}_{N}={\hat{p}}_{2}^{2}+\sum_{\ell}^{2N}\sum_{s}^{{\rm dim}\hat{\mathcal{P}}_{\!\ell}}A_{\ell}^{(s)}\hat{\mathcal{P}}_{\ell}^{(s)},\qquad{\hat{I}}_{N}^{\prime}={\hat{p}}_{1}^{2}+\sum_{\ell}^{2N}\sum_{s}^{{\rm dim}\hat{\mathcal{P}}_{\!\ell}}B_{\ell}^{(s)}\hat{\mathcal{P}}_{\ell}^{(s)},$ (2.6)
where $A_{\ell}^{(s)}$ and $B_{\ell}^{(s)}$ are certain coefficients to be determined. These operators provide quantum symmetries for the Hamiltonian ^$\hat{\mathcal{H}}_{N}$ (2.3) by imposing the condition
$\bigl{[}{\hat{I}}_{N},\hat{\mathcal{H}}_{N}\bigr{]}=\bigl{[}{\hat{I}}_{N}^{\prime},\hat{\mathcal{H}}_{N}\bigr{]}=0,$ (2.7)
which determines the explicit form for the monomials $\hat{\mathcal{P}}_{\ell}^{(s)}$ together with the values of $A_{\ell}^{(s)}$ and $B_{\ell}^{(s)}$. Provided that these quantum symmetries have been found, it follows that the sets (^$(\hat{\mathcal{H}}_{N},\hat{\mathcal{C}},{\hat{I}}_{N})$ and (^$(\hat{\mathcal{H}}_{N},\hat{\mathcal{C}},{\hat{I}}^{\prime}_{N})$ are formed each by three algebraically independent operators, so that ^$\hat{\mathcal{H}}_{N}$ is a quantum superintegrable Hamiltonian. As happens in the classical case (see [3]), the four operators (^$(\hat{\mathcal{H}}_{N},\hat{\mathcal{C}},{\hat{I}}_{N},{\hat{I}}^{\prime}_{N})$ satisfy an algebraic dependence relation.
The second step is to compute the corresponding spectrum, which must exhibit a maximal degeneracy of the energy levels, by means of an algebraic procedure. This requires the obtainment of the associated deformed oscillator algebra [15, 16, 17, 18, 19, 20, 21, 22, 23]; for the use of deformed oscillator algebras in superintegrability and their connection with the symmetry algebra we refer to [24, 25, 26, 27, 28] and references therein. In our case, for a given $N$, this step can be systematized as follows:
  • Define the following operators from the quantum symmetries $(\hat{\mathcal{C}},{\hat{I}}_{N},{\hat{I}}^{\prime}_{N})$:
    $\hat{\mathcal{K}}_{1}:=\hat{\mathcal{C}},\qquad\hat{\mathcal{K}}_{2}:=\frac{1}{2}({\hat{I}}_{N}^{\prime}-{\hat{I}}_{N}),\qquad\hat{\mathcal{K}}_{3}:=[\hat{\mathcal{K}}_{1},\hat{\mathcal{K}}_{2}],$ (2.8)
    which close on a polynomial deformation of $\mathfrak{sl}(2,\mathbb{R})\simeq\mathfrak{so}(2,1)$.
  • Introduce number and ladder (raising/lowering) operators $(\hat{\mathcal{K}},\hat{\mathcal{K}}_{+},\hat{\mathcal{K}}_{-})$ from $(\hat{\mathcal{K}}_{1},\hat{\mathcal{K}}_{2},\hat{\mathcal{K}}_{3})$ via a nonlinear change of basis (which depends on the chosen $N$ through the parameters $\gamma_{k}$), such that
    $\bigl{[}\hat{\mathcal{K}},\hat{\mathcal{K}}_{\pm}\bigr{]}=\pm\hat{\mathcal{K}}_{\pm},\qquad\bigl{[}\hat{\mathcal{K}}_{-},\hat{\mathcal{K}}_{+}\bigr{]}=\Phi\bigl{(}\hat{\mathcal{H}}_{N},\hat{\mathcal{K}}+{\rm Id}\bigr{)}-\Phi\bigl{(}\hat{\mathcal{H}}_{N},\hat{\mathcal{K}}\bigr{)},$ (2.9)
    where ${\rm Id}$ is the identity operator and $\Phi$ is a structure function. The latter is assumed to factorize as
    $\hat{\mathcal{K}}_{+}\hat{\mathcal{K}}_{-}\equiv\Phi\bigl{(}\hat{\mathcal{H}}_{N},\hat{\mathcal{K}}\bigr{)}=\Phi_{1}\bigl{(}\hat{\mathcal{H}}_{N},\hat{\mathcal{K}}\bigr{)}\Phi_{2}\bigl{(}\hat{\mathcal{H}}_{N},\hat{\mathcal{K}}\bigr{)}.$ (2.10)
  • The deformed oscillator algebra arises in a basis $({\hat{B}},{\hat{b}},{\hat{b}}^{\dagger})$ by setting
    ${\hat{B}}:=\hat{\mathcal{K}}-u\,{\rm Id},\qquad{\hat{b}}:=\hat{\mathcal{K}}_{-},\qquad{\hat{b}}^{\dagger}:=\hat{\mathcal{K}}_{+},$ (2.11)
    where $u$ is a constant. The commutation relations read as
    $\bigl{[}{\hat{B}},{\hat{b}}^{\dagger}\bigr{]}={\hat{b}}^{\dagger},\quad\ \bigl{[}{\hat{B}},{\hat{b}}\bigr{]}=-{\hat{b}},\quad\ \bigl{[}{\hat{b}},{\hat{b}}^{\dagger}\bigr{]}=\Phi\bigl{(}\hat{\mathcal{H}}_{N},{\hat{B}}+(u+1){\rm Id}\bigr{)}-\Phi\bigl{(}\hat{\mathcal{H}}_{N},{\hat{B}}+u\,{\rm Id}\bigr{)},$ (2.12)
    where ${\hat{b}}^{\dagger}{\hat{b}}=\Phi\bigl{(}\hat{\mathcal{H}}_{N},{\hat{B}}+u\,{\rm Id}\bigr{)}$.
  • Finally, using the eigenvalues $E$ of ^$\hat{\mathcal{H}}_{N}$ and ${B}$ of ${\hat{B}}$, we consider a finite-dimensional representation of (2.12), that is,
    $\Phi=\Phi({B},E,u)=\Phi_{1}({B},E,u)\Phi_{2}({B},E,u),$ (2.13)
    subjected to the conditions
    $\Phi(0,E,u)=0,\qquad\Phi({n}+1,E,u)=0$ (2.14)
    for a natural number ${n}\in\{1,2,\dots\}$. This set of algebraic equations provides us with constraints on the spectrum and the representation-dependent constant $u$. Hence the solutions take the form $u=u({n})$ and $E=E({n})$, thus leading to $\Phi({B},E({n}),u({n}))\equiv\Phi({B},{n})$ for ${B}\in\{1,\dots,{n}\}$. If $\Phi(B,{n})>0$, the finite-dimensional representation is also unitary. The parameter ${n}$, which characterizes the dimension of the subspace, i.e. of dimension ${n}+1$, determines the principal quantum number $\mathfrak{n}$.
In the following sections 35, we work out the quadratic, cubic and quartic generalized quantum Zernike Hamiltonians ^$\hat{\mathcal{H}}_{N}$ (2.3), corresponding to the values $N=2,3,4$, respectively. For each case, we obtain the quantum symmetries (2.7), the polynomial algebras (2.8) and (2.9), together with the deformed oscillator algebra (2.12). From the latter, the solutions for the possible spectra determined by the equations (2.14) are deduced and analyzed. Furthermore, such new results are also interpreted as superintegrable perturbations of the isotropic oscillator on the two-dimensional sphere, hyperbolic and Euclidean spaces.
While we have not yet fully derived the general expressions for the deformed oscillator algebra (2.12) along with the corresponding spectra for arbitrary $N$, we present two conjectures in section 6 that address these problems. Additionally, we provide a brief overview of the case $N=5$ to further verify the correctness of the results. These conjectures serve as a foundation for future work to establish their proofs.
3 The proper (quadratic) quantum Zernike Hamiltonian
Let us consider the quadratic case of ^$\hat{\mathcal{H}}_{N}$ (2.3) with $N=2$ and arbitrary parameters $\gamma_{1}$ and $\gamma_{2}$. The second-order in the momenta quantum symmetries (2.6) commuting with ^$\hat{\mathcal{H}}_{2}$ (2.7) are found to be
$\begin{split}{\hat{I}}_{2}&={\hat{p}}_{2}^{2}+\gamma_{1}{\hat{q}}_{2}{\hat{p}}_{2}-\gamma_{2}\bigl{(}{\hat{q}}_{2}^{2}({\hat{p}}_{1}^{2}-{\hat{p}}_{2}^{2})-2{\hat{q}}_{1}{\hat{q}}_{2}{\hat{p}}_{1}{\hat{p}}_{2}+{\rm i}{\hat{q}}_{1}{\hat{p}}_{1}+{\rm i}{\hat{q}}_{2}{\hat{p}}_{2}\bigr{)}\\ &={\hat{p}}_{2}^{2}+\gamma_{1}{\hat{q}}_{2}{\hat{p}}_{2}+\gamma_{2}\bigl{(}({\hat{q}}_{1}^{2}+{\hat{q}}_{2}^{2}){\hat{p}}_{2}^{2}-\hat{\mathcal{C}}^{2}\bigr{)},\\[2.0pt] {\hat{I}}^{\prime}_{2}&={\hat{p}}_{1}^{2}+\gamma_{1}{\hat{q}}_{1}{\hat{p}}_{1}+\gamma_{2}\bigl{(}{\hat{q}}_{1}^{2}+{\hat{q}}_{2}^{2}\bigr{)}{\hat{p}}_{1}^{2},\\ \end{split}$ (3.1)
where $\hat{\mathcal{C}}$ is the quantum angular momentum (2.5). These satisfy the relation
^$\hat{\mathcal{H}}_{2}={\hat{I}}_{2}+{\hat{I}}_{2}^{\prime}.$ (3.2)
From $(\hat{\mathcal{C}},{\hat{I}}_{2},{\hat{I}}^{\prime}_{2})$ we introduce the operators $(\hat{\mathcal{K}}_{1},\hat{\mathcal{K}}_{2},\hat{\mathcal{K}}_{3})$ defined in (2.8), which obey the following polynomial commutation relations:
$\begin{split}\bigl{[}\hat{\mathcal{K}}_{1},\hat{\mathcal{K}}_{2}\bigr{]}&=\hat{\mathcal{K}}_{3},\qquad\bigl{[}\hat{\mathcal{K}}_{1},\hat{\mathcal{K}}_{3}\bigr{]}=4\hat{\mathcal{K}}_{2}-2\gamma_{2}\hat{\mathcal{K}}_{1}^{2},\\[2.0pt] \bigl{[}\hat{\mathcal{K}}_{2},\hat{\mathcal{K}}_{3}\bigr{]}&=\bigl{(}\gamma_{1}^{2}+2{\rm i}\gamma_{1}\gamma_{2}+2\gamma_{2}\hat{\mathcal{H}}_{2}\bigr{)}\hat{\mathcal{K}}_{1}+4\gamma_{2}\hat{\mathcal{K}}_{1}\hat{\mathcal{K}}_{2}-2\gamma_{2}\hat{\mathcal{K}}_{3}.\end{split}$ (3.3)
We observe that the corresponding classical functions $I_{2}$, $I_{2}^{\prime}$ and $\mathcal{K}_{i}$’s coming from the quantum operators (3.1) and (2.8) do not exactly coincide with $\mathcal{I}_{2}$, $\mathcal{I}_{2}^{\prime}$ and $\mathcal{L}_{i}$’s as obtained in [3]. Actually, they are related to these through $I_{2}=\mathcal{I}_{2}-\gamma_{2}\mathcal{C}^{2}$, $I_{2}^{\prime}=\mathcal{I}_{2}^{\prime}$, where $\mathcal{C}$ is given in (1.4). This explains why the classical counterpart of the quantum algebra (3.3) does not provide us with the algebra determined by the functions $\mathcal{L}_{i}$’s in [3]. These slight differences are also observable for higher values of $N$, and have their origin in the presence or not of the quantum angular momentum $\hat{\mathcal{C}}$.
We define the number and ladder operators $(\hat{\mathcal{K}},\hat{\mathcal{K}}_{+},\hat{\mathcal{K}}_{-})$ in (2.9) as
$\begin{split}\hat{\mathcal{K}}&:=\frac{1}{2}\hat{\mathcal{K}}_{1},\\ \hat{\mathcal{K}}_{+}&:=\hat{\mathcal{K}}_{2}+\frac{1}{2}\hat{\mathcal{K}}_{3}-\frac{1}{2}\gamma_{2}\hat{\mathcal{K}}_{1}^{2},\\ \hat{\mathcal{K}}_{-}&:=\hat{\mathcal{K}}_{2}-\frac{1}{2}\hat{\mathcal{K}}_{3}-\frac{1}{2}\gamma_{2}\hat{\mathcal{K}}_{1}^{2},\end{split}$ (3.4)
which close on the polynomial symmetry algebra given by
$\bigl{[}\hat{\mathcal{K}},\hat{\mathcal{K}}_{\pm}\bigr{]}=\pm\hat{\mathcal{K}}_{\pm},\qquad\bigl{[}\hat{\mathcal{K}}_{-},\hat{\mathcal{K}}_{+}\bigr{]}=2\bigl{(}\gamma_{1}^{2}+2{\rm i}\gamma_{1}\gamma_{2}+2\gamma_{2}\hat{\mathcal{H}}_{2}\bigr{)}\hat{\mathcal{K}}+16\gamma_{2}^{2}\hat{\mathcal{K}}^{3}.$
The symmetry algebra of the quantum Hamiltonian ^$\hat{\mathcal{H}}_{2}$ becomes, over this basis, a cubic polynomial algebra $\mathfrak{sl}^{(3)}(2,\mathbb{R})$, which is just the so-called Higgs algebra [29] (see also [30, 31, 32, 33, 34] and references therein), as already pointed out in [5].
In order to construct a deformed Heisenberg algebra and compute the corresponding spectrum of ^$\hat{\mathcal{H}}_{2}$, we need to find the structure function $\Phi$ in (2.9), which turns out to be
$\begin{split}\hat{\mathcal{K}}_{+}\hat{\mathcal{K}}_{-}=\Phi(\hat{\mathcal{H}}_{2},\hat{\mathcal{K}})&=\frac{1}{4}\bigl{(}4\gamma_{2}-2{\rm i}\gamma_{1}+\hat{\mathcal{H}}_{2}\bigr{)}\hat{\mathcal{H}}_{2}-\bigl{(}\gamma_{1}^{2}+2{\rm i}\gamma_{1}\gamma_{2}+2\gamma_{2}\hat{\mathcal{H}}_{2}\bigr{)}\hat{\mathcal{K}}\\ &\qquad+\bigl{(}\gamma_{1}^{2}+4\gamma_{2}^{2}+2{\rm i}\gamma_{1}\gamma_{2}+2\gamma_{2}\hat{\mathcal{H}}_{2}\bigr{)}\hat{\mathcal{K}}^{2}-8\gamma_{2}^{2}\hat{\mathcal{K}}^{3}+4\gamma_{2}^{2}\hat{\mathcal{K}}^{4}.\end{split}$
Following (2.10), the factorization of $\Phi$ is given by
$\begin{split}\Phi(\hat{\mathcal{H}}_{2},\hat{\mathcal{K}})&=\Phi_{1}(\hat{\mathcal{H}}_{2},\hat{\mathcal{K}})\Phi_{2}(\hat{\mathcal{H}}_{2},\hat{\mathcal{K}}),\qquad\bigl{[}\Phi_{1}(\hat{\mathcal{H}}_{2},\hat{\mathcal{K}}),\Phi_{2}(\hat{\mathcal{H}}_{2},\hat{\mathcal{K}})\bigr{]}=0,\\[2.0pt] \Phi_{1}(\hat{\mathcal{H}}_{2},\hat{\mathcal{K}})&=\frac{1}{4}\bigl{(}\hat{\mathcal{H}}_{2}-2{\rm i}\gamma_{1}\hat{\mathcal{K}}+4\gamma_{2}\hat{\mathcal{K}}^{2}\bigr{)},\\[2.0pt] \Phi_{2}(\hat{\mathcal{H}}_{2},\hat{\mathcal{K}})&=\hat{\mathcal{H}}_{2}-2({\rm i}\gamma_{1}-2\gamma_{2}){\rm Id}+2({\rm i}\gamma_{1}-4\gamma_{2})\hat{\mathcal{K}}+4\gamma_{2}\hat{\mathcal{K}}^{2}\\ &=\hat{\mathcal{H}}_{2}+2{\rm i}\gamma_{1}\bigl{(}\hat{\mathcal{K}}-{\rm Id}\bigr{)}+4\gamma_{2}\bigl{(}\hat{\mathcal{K}}-{\rm Id}\bigr{)}^{2},\end{split}$ (3.5)
which gives rise to the deformed oscillator algebra (2.12) under the change of operators (2.11).
3.1 Spectrum of the quadratic system
At this stage, we have all the information to compute the spectrum $E$ of ^$\hat{\mathcal{H}}_{2}$ in (2.4) algebraically. Using (3.5), it follows that the finite-dimensional representation of (2.12) yields the structure function $\Phi$ (2.13) as
$\begin{split}\Phi({B},E,u)&=\Phi_{1}({B},E,u)\Phi_{2}({B},E,u),\\ \Phi_{1}({B},E,u)&=\frac{1}{4}\bigl{(}E-2{\rm i}\gamma_{1}({B}+u)+4\gamma_{2}({B}+u)^{2}\bigr{)},\\ \Phi_{2}({B},E,u)&=E+2{\rm i}\gamma_{1}({B}+u-1)+4\gamma_{2}({B}+u-1)^{2},\end{split}$ (3.6)
where we recall that $E$ and ${B}$ are the eigenvalues of ^$\hat{\mathcal{H}}_{2}$ and ${\hat{B}}$, respectively. Taking into account the two constraints in (2.14), the result can be summarized in the following
Proposition 1
The set of equations (2.14) yields four types of solutions for the representation-dependent constant $u$ and the spectrum of the Hamiltonian ^$\hat{\mathcal{H}}_{2}$ in (2.4) depending on the parameter ${n}\in\{1,2,\dots\}$, that is, $u=u({n})$ and $E=E({n})$. Introducing these expressions into (3.6), the corresponding structure functions $\Phi({B},E({n}),u({n}))\equiv\Phi({B},{n})$ for ${B}=\{1,\dots,{n}\}$ are obtained. The final result is displayed in Table 1.
Table 1: The four types of solutions of the equations (2.14) with the corresponding representation dependent constant $u({n})$, spectrum $E({n})$ and structure function $\Phi(B,{n})$ (3.6) where ${n}\in\mathbb{N}^{\ast}$ and $B\in\{1,\dots,{n}\}$.
Type I $u_{\rm I}=-\frac{{n}}{2}$   $E_{\rm I}=-{n}({\rm i}\gamma_{1}+\gamma_{2}{n})$
$\Phi_{\rm I}=-{B}({B}-{n}-1)\bigl{(}{\rm i}\gamma_{1}+2\gamma_{2}({B}-1)\bigr{)}\bigl{(}{\rm i}\gamma_{1}-2\gamma_{2}({B}-{n})\bigr{)}$
Type II $u_{\rm II}=-\frac{{n}}{2}$  $E_{\rm II}=({n}+2)\bigr{(}{\rm i}\gamma_{1}-\gamma_{2}({n}+2)\bigl{)}$
$\Phi_{\rm II}=-{B}({B}-{n}-1)\bigl{(}{\rm i}\gamma_{1}-2\gamma_{2}({B}+1)\bigr{)}\bigl{(}{\rm i}\gamma_{1}+2\gamma_{2}({B}-{n}-2)\bigr{)}$
Type III $u_{\rm III}=-\frac{1}{2}\left({n}-1+\frac{{\rm i}\gamma_{1}}{2\gamma_{2}}\right)$ $\ \displaystyle E_{\rm III}=-\frac{\gamma_{1}^{2}}{4\gamma_{2}}-\gamma_{2}({n}+1)^{2}$
$\Phi_{\rm III}={B}({B}-{n}-1)\bigl{(}{\rm i}\gamma_{1}-2\gamma_{2}({B}+1)\bigr{)}\bigl{(}{\rm i}\gamma_{1}-2\gamma_{2}({B}-{n})\bigr{)}$
Type IV $u_{\rm IV}=-\frac{1}{2}\left({n}+1-\frac{{\rm i}\gamma_{1}}{2\gamma_{2}}\right)$ $\ \displaystyle E_{\rm IV}=-\frac{\gamma_{1}^{2}}{4\gamma_{2}}-\gamma_{2}({n}+1)^{2}$
$\Phi_{\rm IV}={B}({B}-{n}-1)\bigl{(}{\rm i}\gamma_{1}+2\gamma_{2}({B}-1)\bigr{)}\bigl{(}{\rm i}\gamma_{1}+2\gamma_{2}({B}-{n}-2)\bigr{)}$
Some remarks are in order. The solutions of type III and IV give rise to the same spectrum, $E_{\rm III}=E_{\rm IV}$, although the structure functions $\Phi$ and hence the deformed oscillator algebras in (2.12) are different. The four types of spectra take real values whenever the parameter $\gamma_{1}$ is a pure imaginary number, keeping $\gamma_{2}$ as a real parameter. This suggests to introduce arbitrary real parameters $\alpha$ and $\beta$ as defined in (1.5), and appearing within the Zernike Hamiltonian (1.2). With these convention, we find that
$E_{\rm I}=-{n}(\beta+\alpha{n}),\qquad E_{\rm II}=({n}+2)\bigr{(}\beta-\alpha({n}+2)\bigl{)},\qquad E_{\rm III}=E_{\rm IV}=\frac{\beta^{2}}{4\alpha}-\alpha({n}+1)^{2}.$ (3.7)
Different possibilities appear depending on the specific values of $\alpha$ and $\beta$. Quite unexpectedly, the quantum Zernike system corresponding to the values $\alpha=-1$ and $\beta=-2$ emerges as a very ‘special’ case, as the four types of solutions merge into the same, namely
$E({n})={n}({n}+2),\qquad u({n})=-\frac{{n}}{2},\qquad\Phi({B},{n})=4{B}^{2}({B}-{n}-1)^{2},$
where ${n}\in\{1,2,\dots\}$ and ${B}\in\{1,\dots,{n}\}$. As $\Phi({B},n)>0$ holds, there is no restriction on the value of ${n}$, and the finite-dimensional representation is always unitary. Identifying the parameter ${n}$ with the principal quantum number $\mathfrak{n}$, we recover the expression
$E=\mathfrak{n}(\mathfrak{n}+2)$ (3.8)
given in [5] (cf. eq. (16)). Observe that, strictly speaking, ${n}\in\{1,2,\dots\}$, while $\mathfrak{n}\in\{0,1,2,\dots\}$, indicating that the difference relies on the ground state.
3.2 Application to the spherical and hyperbolic oscillators
In addition, it is worth remarking that, by means of a canonical transformation, it has been shown in [3] (see also [8]) that the classical Zernike system $\mathcal{H}_{\rm Zk}\equiv\mathcal{H}_{2}$ (2.3) can be regarded as an isotropic oscillator on the two-dimensional sphere $\mathbf{S}^{2}$ and the hyperbolic space $\mathbf{H}^{2}$. Within this interpretation, the imaginary coefficient $\gamma_{1}$ is related to the frequency $\omega$ in the form $\gamma_{1}=2{\rm i}\omega$, while the second one is just minus the constant curvature of the space $\gamma_{2}=-\kappa$. For the sake of completeness we recall the explicit form of $\mathcal{H}_{2}$ as a natural Hamiltonian in terms of a curved kinetic energy $\mathcal{T}_{\kappa}$ and a central potential $\mathcal{U}_{\kappa}\propto\tan^{2}(\sqrt{\kappa}\rho)/\kappa$:
$\mathcal{H}_{2}=\mathcal{T}_{\kappa}+\mathcal{U}_{\kappa}(\rho),\qquad\mathcal{T}_{\kappa}=p_{\rho}^{2}+\frac{\kappa\,p_{\phi}^{2}}{\sin^{2}(\sqrt{\kappa}\rho)},\qquad\mathcal{U}_{\kappa}(\rho)=-\frac{\gamma_{1}^{2}}{4\kappa}\tan^{2}(\sqrt{\kappa}\rho),$ (3.9)
where $\rho$ is a geodesic radial coordinate, i.e., the distance between the particle and the origin on the space measured along the geodesic joining both points, while $\phi\in[0,2\pi)$ is an ordinary angle. Therefore, $\mathcal{H}_{2}$ covers the well-known two-dimensional isotropic spherical or Higgs oscillator on $\mathbf{S}^{2}$ ($\gamma_{2}<0$, $\kappa>0$), the hyperbolic system on $\mathbf{H}^{2}$ ($\gamma_{2}>0$, $\kappa<0$) and the usual one on the Euclidean plane $\mathbf{E}^{2}$ $(\gamma_{2}=\kappa=0)$ (see [35, 36, 37, 38, 39] and references therein). For the ‘standard’ values of the curvature $\kappa\in\{+1,0,-1\}$ we find that (3.9) gives
$\begin{array}[]{lll}\mathbf{S}^{2}&(\kappa=+1)\!\!:&\quad\displaystyle\mathcal{H}_{2}=p_{\rho}^{2}+\frac{\,p_{\phi}^{2}}{\sin^{2}\!\rho}-\frac{\gamma_{1}^{2}}{4}\tan^{2}\!\rho.\\[10.0pt] \mathbf{E}^{2}&(\kappa=0)\!\!:&\quad\displaystyle\mathcal{H}_{2}=p_{\rho}^{2}+\frac{\,p_{\phi}^{2}}{\rho^{2}}-\frac{\gamma_{1}^{2}}{4}\,\rho^{2}.\\[10.0pt] \mathbf{H}^{2}&(\kappa=-1)\!\!:&\quad\displaystyle\mathcal{H}_{2}=p_{\rho}^{2}+\frac{\,p_{\phi}^{2}}{\sinh^{2}\!\rho}-\frac{\gamma_{1}^{2}}{4}\tanh^{2}\!\rho.\end{array}$
From this perspective, the solutions of type III and IV in table 1 are not well-defined for the flat Euclidean case with $\kappa=\gamma_{2}\to 0$.
Although both quantum curved spherical and hyperbolic oscillators have already been fully solved in [35, 38] obtaining the corresponding spectrum and eigenfunctions, our algebraic procedure allows us to recover their spectrum straightforwardly, that shows deep differences between the spherical and hyperbolic cases, but both reproduce the Euclidean oscillator under the flat limit $\kappa\to 0$.
Let us focus on the solution of type I in table 1, set $\gamma_{1}=2{\rm i}$ i.e. $\beta=-2$ (this means to fix $\omega=1$) and express $\gamma_{2}=\alpha=-\kappa$, allowing any value for the curvature $\kappa$. We analyze separately the three particular systems according to $\kappa$ and for simplicity hereafter we drop the index ‘I’.
The case with $\kappa=0$ leads to the isotropic oscillator on $\mathbf{E}^{2}$ such that
$E(n)=2n,\qquad\Phi=4{B}({n}+1-{B}),$ (3.10)
with ${n}\in\{1,2,\dots\}$, ${B}=\{1,\dots,{n}\}$ so that $\Phi>0$. Hence, as expected, the spectrum is linear in ${n}$, unbounded and with a constant difference $E(n+1)-E(n)=2$. This implies that there exists an infinite set of bound states.
The spherical or Higgs oscillator arises on $\mathbf{S}^{2}$ when $\kappa>0$ leading to the expressions
$E(n)=2n+\kappa n^{2},\qquad\Phi=4{B}({n}+1-{B})\bigr{(}1+\kappa(B-1)\bigl{)}\bigr{(}1+\kappa(n-B)\bigl{)},$ (3.11)
where again ${n}\in\{1,2,\dots\}$, ${B}=\{1,\dots,{n}\}$ and $\Phi>0$. Now the spectrum is quadratic in ${n}$, with higher values than the Euclidean oscillator, and unbounded. Nevertheless, the difference of the energy between two consecutive states is no longer constant but increases linearly with ${n}$ as
$E(n+1)-E(n)=2+\kappa(2n+1).$ (3.12)
Consequently, there also exists an infinite number of bound states. We represent the first energies (3.11) for three values of the curvature and those of the Euclidean oscillator (3.10) in figure 1. Obviously, as the curvature grows the differences between the spherical and Euclidean spectra become larger.
For the hyperbolic oscillator on $\mathbf{H}^{2}$ we write $\kappa=-|\kappa|$, obtaining that
$E(n)=2n-|\kappa|n^{2},\qquad\Phi=4{B}({n}+1-{B})\bigr{(}1-|\kappa|(B-1)\bigl{)}\bigr{(}1-|\kappa|(n-B)\bigl{)}.$ (3.13)
In contrast to (3.12), the difference of the energies is decreasing and turns into
$E(n+1)-E(n)=2-|\kappa|(2n+1).$ (3.14)
In order to relate these results to those previously deduced in [35, 38] and smoothly recover the Euclidean system, we require a positive energy spectrum $E(n)$ (3.13) and positive differences of two consecutive energies (3.14). The first condition leads to $2>|\kappa|n$, thus the lower value $n=1$ determines a limiting value for the curvature: $|\kappa|<2$. The second requirement implies that there exists a maximum value $n_{\rm max}$ for the positive integer $n\geq 1$, which is the greatest integer number fulfilling the constraint obtained by setting $n_{\rm max}=n+1\geq 2$ in (3.14), namely
$2\leq n_{\rm max}<\frac{1}{|\kappa|}+\frac{1}{2}.$
The single state (ground case) for the extreme value $n=n_{\rm max}=1$ can be included into a unique condition, such that both restrictions can be written as
$21\leq n_{\rm max}<\frac{1}{|\kappa|}+\frac{1}{2}\quad\Longrightarrow\quad|\kappa|<$ 2 . (3.15)
The implication is a direct consequence of taking the lowest $n_{\rm max}=1$. Moreover, observe that the specific values appearing in these relations depend not only on the curvature but also on the particular choice of the parameter $\beta=-2$ (hence on the frequency $\omega$). And, clearly, if $|\kappa|\to 0$, then $n_{\rm max}\to\infty$, recovering the Euclidean case. Provided that the value of $n_{\rm max}$ has been fixed, the relation (3.15) gives
$\begin{array}[]{ll}n_{\rm max}=1\!\!:&\quad\displaystyle\frac{2}{3}\leq|\kappa|<2.\\[10.0pt] n_{\rm max}>1\!\!:&\quad\displaystyle\frac{2}{2n_{\rm max}+1}\leq|\kappa|<\frac{2}{2n_{\rm max}-1},\end{array}$ (3.16)
that is, $32/5\leq|\kappa|<2/$ 3 for $n_{\rm max}=2$, $52/7\leq|\kappa|<2/$ 5 for $n_{\rm max}=3$, and so on. Thus one can look for a certain number of states choosing a value for $n_{\rm max}$ and obtain all possible compatible values for the curvature.
Let us now prove that the restriction (3.15) always ensure that $\Phi>0$ in (3.13) showing that $\bigr{(}1-|\kappa|(B-1)\bigl{)}>0$ and $\bigr{(}1-|\kappa|(n-B)\bigl{)}>0$. Taking into account that ${B}=\{1,\dots,{n}\}$ the maximum value for $B\geq 1$ is just $n_{\rm max}$ (3.15). For the first factor we find that
$\frac{1}{|\kappa|}+1>n_{\rm max}\geq B\geq 1.$
And for the second factor we take the extreme values $n=n_{\rm max}$ and $B=1$:
$\frac{1}{|\kappa|}>n-B\quad\Longleftrightarrow\quad\frac{1}{|\kappa|}>n_{\rm max}-1.$
Therefore and in agreement with [35, 38], the isotropic hyperbolic oscillator has a positive finite discrete spectrum with ${n}\in\{1,\dots,n_{\rm max}\}$ and there exist unbounded states which may be provided by one-dimensional hyperbolic Pöschl–Teller potentials, as pointed out in Theorem 6 in [38]. We illustrate these results in figure 2, plotting the discrete spectra for three values of $n_{\rm max}$ (so choosing a value for the curvature) which are always lower than the Euclidean ones (3.10).
[…]
4 The cubic quantum Zernike Hamiltonian
In this section we study the quantum symmetries and deduce the possible spectra for the cubic Hamiltonian ^$\hat{\mathcal{H}}_{3}$ (2.3) depending on three arbitrary parameters $\gamma_{1}$, $\gamma_{2}$ and $\gamma_{3}$. It is worthy to be observed that the cubic $\gamma_{3}$-term in ^$\hat{\mathcal{H}}_{3}$ can be interpreted as a superintegrable perturbation of the quantum Zernike Hamiltonian ^$\hat{\mathcal{H}}_{\rm Zk}$ in (2.2). Indeed, we require that all the expressions for ^$\hat{\mathcal{H}}_{2}$ in (2.3) obtained in the previous section are recovered whenever $\gamma_{3}$ is put equal to zero.
The two quantum symmetries (2.6) of ^$\hat{\mathcal{H}}_{3}$, that are of third-order in the momenta, are obtained after some computation, and adopt the form
$\begin{split}{\hat{I}}_{3}&={\hat{p}}_{2}^{2}+\gamma_{1}{\hat{q}}_{2}{\hat{p}}_{2}+\gamma_{2}\bigl{(}({\hat{q}}_{1}^{2}+{\hat{q}}_{2}^{2}){\hat{p}}_{2}^{2}-\hat{\mathcal{C}}^{2}\bigr{)}\\[2.0pt] &\qquad+\gamma_{3}\bigl{(}{\hat{q}}_{2}^{3}({\hat{p}}_{2}^{3}-{\hat{p}}_{1}^{2}{\hat{p}}_{2})+({\hat{q}}_{1}^{3}+3{\hat{q}}_{1}{\hat{q}}_{2}^{2}){\hat{p}}_{1}{\hat{p}}_{2}^{2}-3{\rm i}{\hat{q}}_{2}^{2}{\hat{p}}_{2}^{2}-3{\rm i}{\hat{q}}_{1}{\hat{q}}_{2}{\hat{p}}_{1}{\hat{p}}_{2}-{\hat{q}}_{2}{\hat{p}}_{2}\bigr{)},\\[4.0pt] {\hat{I}}^{\prime}_{3}&={\hat{p}}_{1}^{2}+\gamma_{1}{\hat{q}}_{1}{\hat{p}}_{1}+\gamma_{2}\bigl{(}{\hat{q}}_{1}^{2}+{\hat{q}}_{2}^{2}\bigr{)}{\hat{p}}_{1}^{2}\\[2.0pt] &\qquad+\gamma_{3}\bigl{(}{\hat{q}}_{1}^{3}({\hat{p}}_{1}^{3}-{\hat{p}}_{1}{\hat{p}}_{2}^{2})+({\hat{q}}_{2}^{3}+3{\hat{q}}_{1}^{2}{\hat{q}}_{2}){\hat{p}}_{1}^{2}{\hat{p}}_{2}-3{\rm i}{\hat{q}}_{1}^{2}{\hat{p}}_{1}^{2}-3{\rm i}{\hat{q}}_{1}{\hat{q}}_{2}{\hat{p}}_{1}{\hat{p}}_{2}-{\hat{q}}_{1}{\hat{p}}_{1}\bigr{)}.\end{split}$
They satisfy the relation (3.2). The operators $(\hat{\mathcal{K}}_{1},\hat{\mathcal{K}}_{2},\hat{\mathcal{K}}_{3})$, as defined in (2.8), close on the polynomial algebra given by
$\begin{split}\bigl{[}\hat{\mathcal{K}}_{1},\hat{\mathcal{K}}_{2}\bigr{]}&=\hat{\mathcal{K}}_{3},\qquad\bigl{[}\hat{\mathcal{K}}_{1},\hat{\mathcal{K}}_{3}\bigr{]}=4\hat{\mathcal{K}}_{2}-2\gamma_{2}\hat{\mathcal{K}}_{1}^{2},\\[2.0pt] \bigl{[}\hat{\mathcal{K}}_{2},\hat{\mathcal{K}}_{3}\bigr{]}&=\bigl{(}\gamma_{1}^{2}+2{\rm i}\gamma_{1}\gamma_{2}-4\gamma_{1}\gamma_{3}+2(\gamma_{2}+3{\rm i}\gamma_{3})\hat{\mathcal{H}}_{3}\bigr{)}\hat{\mathcal{K}}_{1}+4\gamma_{2}\hat{\mathcal{K}}_{1}\hat{\mathcal{K}}_{2}-2\gamma_{2}\hat{\mathcal{K}}_{3}\\ &\qquad-4\gamma_{3}(\gamma_{1}-{\rm i}\gamma_{2}-\gamma_{3})\hat{\mathcal{K}}_{1}^{3}+3\gamma_{3}^{2}\hat{\mathcal{K}}_{1}^{5}.\end{split}$
The number and ladder operators $(\hat{\mathcal{K}},\hat{\mathcal{K}}_{+},\hat{\mathcal{K}}_{-})$ are defined exactly as in (3.4), and satisfy the commutation relations (2.9) with
$\begin{split}\bigl{[}\hat{\mathcal{K}}_{-},\hat{\mathcal{K}}_{+}\bigr{]}&=2\bigl{(}\gamma_{1}^{2}+2{\rm i}\gamma_{1}\gamma_{2}-4\gamma_{1}\gamma_{3}+2(\gamma_{2}+3{\rm i}\gamma_{3})\hat{\mathcal{H}}_{3}\bigr{)}\hat{\mathcal{K}}\\ &\qquad+16\bigl{(}\gamma_{2}^{2}+2\gamma_{3}^{2}-2\gamma_{1}\gamma_{3}+2{\rm i}\gamma_{2}\gamma_{3}\bigr{)}\hat{\mathcal{K}}^{3}+96\gamma_{3}^{2}\hat{\mathcal{K}}^{5}.\end{split}$
The symmetry algebra of the cubic quantum Hamiltonian ^$\hat{\mathcal{H}}_{3}$ hence corresponds to a fifth-order polynomial Higgs-type algebra $\mathfrak{sl}^{(5)}(2,\mathbb{R})$ (see [3] for the classical picture).
On the other hand, the structure function $\Phi$ (2.9) is determined by
$\begin{split}\hat{\mathcal{K}}_{+}\hat{\mathcal{K}}_{-}&=\Phi(\hat{\mathcal{H}}_{3},\hat{\mathcal{K}})\\ &=\frac{1}{4}\bigl{(}4\gamma_{2}-2{\rm i}\gamma_{1}+8{\rm i}\gamma_{3}+\hat{\mathcal{H}}_{3}\bigr{)}\hat{\mathcal{H}}_{3}-\bigl{(}\gamma_{1}^{2}+2{\rm i}\gamma_{1}\gamma_{2}-4\gamma_{1}\gamma_{3}+2(\gamma_{2}+3{\rm i}\gamma_{3})\hat{\mathcal{H}}_{3}\bigr{)}\hat{\mathcal{K}}\\ &\qquad+\bigl{(}\gamma_{1}^{2}+4\gamma_{2}^{2}+2{\rm i}\gamma_{1}\gamma_{2}-12\gamma_{1}\gamma_{3}+8{\rm i}\gamma_{2}\gamma_{3}+2(\gamma_{2}+3{\rm i}\gamma_{3})\hat{\mathcal{H}}_{3}\bigr{)}\hat{\mathcal{K}}^{2}\\[2.0pt] &\qquad-8\bigl{(}\gamma_{2}^{2}+2\gamma_{3}^{2}-2\gamma_{1}\gamma_{3}+2{\rm i}\gamma_{2}\gamma_{3}\bigr{)}\hat{\mathcal{K}}^{3}+4\bigl{(}\gamma_{2}^{2}+12\gamma_{3}^{2}-2\gamma_{1}\gamma_{3}+2{\rm i}\gamma_{2}\gamma_{3}\bigr{)}\hat{\mathcal{K}}^{4}\\[2.0pt] &\qquad-48\gamma_{3}^{2}\hat{\mathcal{K}}^{5}+16\gamma_{3}^{2}\hat{\mathcal{K}}^{6},\end{split}$
which is then factorized as (see (2.10))
$\begin{split}\Phi(\hat{\mathcal{H}}_{3},\hat{\mathcal{K}})&=\Phi_{1}(\hat{\mathcal{H}}_{3},\hat{\mathcal{K}})\Phi_{2}(\hat{\mathcal{H}}_{3},\hat{\mathcal{K}}),\qquad\bigl{[}\Phi_{1}(\hat{\mathcal{H}}_{3},\hat{\mathcal{K}}),\Phi_{2}(\hat{\mathcal{H}}_{3},\hat{\mathcal{K}})\bigr{]}=0,\\ \Phi_{1}(\hat{\mathcal{H}}_{3},\hat{\mathcal{K}})&=\frac{1}{4}\bigl{(}\hat{\mathcal{H}}_{3}-2{\rm i}\gamma_{1}\hat{\mathcal{K}}+4\gamma_{2}\hat{\mathcal{K}}^{2}+8{\rm i}\gamma_{3}\hat{\mathcal{K}}^{3}\bigr{)},\\[2.0pt] \Phi_{2}(\hat{\mathcal{H}}_{3},\hat{\mathcal{K}})&=\hat{\mathcal{H}}_{3}-2({\rm i}\gamma_{1}-2\gamma_{2}-4{\rm i}\gamma_{3}){\rm Id}+2({\rm i}\gamma_{1}-4\gamma_{2}-12{\rm i}\gamma_{3})\hat{\mathcal{K}}\\ &\qquad+4(\gamma_{2}+6{\rm i}\gamma_{3})\hat{\mathcal{K}}^{2}-8{\rm i}\gamma_{3}\hat{\mathcal{K}}^{3}\\ &=\hat{\mathcal{H}}_{3}+2{\rm i}\gamma_{1}\bigl{(}\hat{\mathcal{K}}-{\rm Id}\bigr{)}+4\gamma_{2}\bigl{(}\hat{\mathcal{K}}-{\rm Id}\bigr{)}^{2}-8{\rm i}\gamma_{3}\bigl{(}\hat{\mathcal{K}}-{\rm Id}\bigr{)}^{3}.\end{split}$
It is routine to verify that with the change of operators (2.11), we get the deformed oscillator algebra (2.12).
4.1 Spectrum of the cubic system
In terms of the finite-dimensional representation of the deformed oscillator algebra (2.12), the structure function $\Phi$ of (2.13) adopts the expression
$\begin{split}\Phi({B},E,u)&=\Phi_{1}({B},E,u)\Phi_{2}({B},E,u),\\ \Phi_{1}({B},E,u)&=\frac{1}{4}\bigl{(}E-2{\rm i}\gamma_{1}({B}+u)+4\gamma_{2}({B}+u)^{2}+8{\rm i}\gamma_{3}({B}+u)^{3}\bigr{)},\\ \Phi_{2}({B},E,u)&=E+2{\rm i}\gamma_{1}({B}+u-1)+4\gamma_{2}({B}+u-1)^{2}-8{\rm i}\gamma_{3}({B}+u-1)^{3}.\end{split}$ (4.1)
The solution can thus be summarized in the following statement.
Proposition 2
(i) The set of equations (2.14) coming from (4.1) yields ten types of solutions for the representation-dependent constant $u=u({n})$ and the spectrum $E=E({n})$ of the Hamiltonian ^$\hat{\mathcal{H}}_{3}$ in (2.4) depending on the parameter ${n}\in\{1,2,\dots\}$. They lead to the structure function $\Phi({B},E({n}),u({n}))\equiv\Phi({B},{n})$ for ${B}=\{1,\dots,{n}\}$ via the factorization (4.1).
(ii) The requirement that the solutions are well-defined for any values of $\gamma_{1}$ and $\gamma_{2}$ for $\gamma_{3}=0$ reduces the ten types of solutions to only two, given by
[…]
To illustrate the situation with the seven discarded solutions for the limit $\gamma_{3}\to 0$, let us consider, for instance, one of them given by
$u=-\frac{{n}}{2}-\frac{1}{2\gamma_{3}}\sqrt{\gamma_{3}\bigl{(}\gamma_{1}-2{\rm i}\gamma_{2}{n}-3\gamma_{3}{n}^{2}\bigr{)}},\qquad E=-\frac{1}{\gamma_{3}}\bigl{(}\gamma_{2}-2{\rm i}\gamma_{3}{n}\bigr{)}\bigl{(}\gamma_{1}-2{\rm i}\gamma_{2}{n}-4\gamma_{3}{n}^{2}\bigr{)}.$ (4.2)
For $\gamma_{1}\neq 0$ or $\gamma_{2}\neq 0$, it is easily seen that $u$, $E$ or both are not well-defined for $\gamma_{3}=0$.
The two consistent types of solutions given in proposition 2 constitute a cubic extension of those presented in proposition 1 and displayed in Table 1. We observe that types III and IV do not admit a cubic generalization, implying that they are specific solutions for the quadratic Hamiltonian (2.3) with $N=2$. The spectrum corresponding to types I and II takes real values whenever the coefficients $\gamma_{1}$ and $\gamma_{3}$ are pure imaginary numbers, while $\gamma_{2}$ is real. By considering (1.5), as before, we denote
$\gamma_{1}=-\rm i\beta,\qquad\gamma_{2}=\alpha,\qquad\gamma_{3}={\rm i}\mu,\qquad\{\alpha,\beta,\mu\}\in\mathbb{R},$ (4.3)
finding that
$\begin{split}E_{\rm I}&=-\bigl{(}\beta{n}+\alpha{n}^{2}+\mu{n}^{3}\bigr{)},\\ E_{\rm II}&=\beta({n}+2)-\alpha({n}+2)^{2}+\mu({n}+2)^{3},\end{split}$ (4.4)
to be compared with (3.7). Furthermore, if we set $\alpha=-1$ and $\beta=-2$ and identify ${n}$ with the principal quantum number $\mathfrak{n}$, we obtain two possible superintegrable perturbations of the spectrum (3.8) of the Zernike Hamiltonian ^$\hat{\mathcal{H}}_{\rm Zk}$ in (1.1), reading as
$E_{\rm I}=\mathfrak{n}(\mathfrak{n}+2)-\mu\mathfrak{n}^{3},\qquad E_{\rm II}=\mathfrak{n}(\mathfrak{n}+2)+\mu(\mathfrak{n}+2)^{3}.$
4.2 Perturbations of the Euclidean, spherical and hyperbolic oscillators
Following the analysis in section 3.2, we can interpret the results of proposition 2 as superintegrable cubic perturbations of the isotropic oscillator on $\mathbf{E}^{2}$, $\mathbf{S}^{2}$ and $\mathbf{H}^{2}$. To this end, let us consider the solution of type I (4.4) (eliminating the index ‘I’), again set $\beta=-2$ and $\alpha=-\kappa$, but now keeping an arbitrary real parameter $\mu$, which characterizes the perturbation. Thus, we have six possibilities, according to the curvature $\kappa$ and the two signs of $\mu$.
For the flat Euclidean oscillator with $\kappa=0$ we start with a negative value for the perturbation parameter $\mu=-|\mu|$ obtaining that
$E(n)=2n+|\mu|n^{3},\qquad E(n+1)-E(n)=2+|\mu|(3n^{2}+3n+1).$ (4.5)
Hence, the spectrum is cubic, unbounded and the energy between two consecutive states grows quadratically with ${n}$. It is worth noting that the behaviour of the spectrum (4.5) is quite similar to that of the spherical or Higgs oscillator (3.11) and (3.12), so we can say that it is a cubic spherical perturbation of the Euclidean oscillator. In figure 3 we plot the first energies (4.5) for three negative values of $\mu$ together with those of the non-perturbed Euclidean oscillator (3.10) (to be compared with figure 1).
Then, we take a positive value for $\mu$ on $\mathbf{E}^{2}$ finding that
$E(n)=2n-\mu n^{3},\qquad E(n+1)-E(n)=2-\mu(3n^{2}+3n+1).$ (4.6)
We require a positive spectrum and positive differences of two consecutive energies. As for the hyperbolic oscillator (see (3.15) and (3.16)) such conditions yield a maximum value $n_{\rm max}$ for the positive integer $n\geq 1$ which is the greatest integer number that satisfies the following constraint
$21\leq n_{\rm max}<\frac{1}{2}+\frac{1}{6}\sqrt{\frac{24}{\mu}-3}\quad\Longrightarrow\quad\mu<$ 2 . (4.7)
Here, the perturbation parameter can be expressed in terms of a fixed value of $n_{\rm max}$ in the form
$\begin{array}[]{ll}n_{\rm max}=1\!\!:&\quad\displaystyle\frac{2}{7}\leq\mu<2.\\[10.0pt] n_{\rm max}>1\!\!:&\quad\displaystyle\frac{2}{1+3n_{\rm max}(n_{\rm max}+1)}\leq\mu<\frac{2}{1+3n_{\rm max}(n_{\rm max}-1)}.\end{array}$
In this respect, we can say that (4.6) and (4.7) determine a cubic hyperbolic perturbation of the Euclidean oscillator that leads to a positive finite discrete spectrum with ${n}\in\{1,\dots,n_{\rm max}\}$. We represent the discrete spectrum for three values of $n_{\rm max}$ (so for $\mu$) in figure 4, which are lower than the Euclidean ones (3.10), and worthy of comparison with figure 2.
Figure 3: The discrete spectrum (4.5) of the cubic spherical perturbation of the oscillator on $\mathbf{E}^{2}$ for the first ten states $1\leq n\leq 10$ according to three values of the parameter $\mu\in\{-0.05,\,-0.025,\,-0.01\}$ starting from the upper dots together with the Euclidean spectrum (3.10) with $\mu=0$ in the lowest values.
[…]
Now we consider the sphere $\mathbf{S}^{2}$ with positive curvature $\kappa>0$ and a negative parameter $\mu=-|\mu|$. Thus we arrive at
$E(n)=2n+\kappa n^{2}+|\mu|n^{3},\qquad E(n+1)-E(n)=2+\kappa(2n+1)+|\mu|(3n^{2}+3n+1).$
As a consequence, the spectrum is discrete and unbounded, but the presence of $\mu<0$ entails a higher energy growth than in the spherical oscillator (3.11) and (3.12), such that it can be regarded as a cubic spherical perturbation of the spherical-Higgs oscillator.
The above situation changes drastically when $\mu>0$. Actually, in this case
$E(n)=2n+\kappa n^{2}-\mu n^{3},\qquad E(n+1)-E(n)=2+\kappa(2n+1)-\mu(3n^{2}+3n+1),$ (4.8)
and the requirements for getting a positive spectrum and positive differences of two consecutive energies imply that there is a maximum value $n_{\rm max}$ for the positive integer $n\geq 1$ that has to satisfy the inequality
$1\leq n_{\rm max}<\frac{1}{2}+\frac{\kappa}{3\mu}+\frac{1}{6}\sqrt{\frac{4\kappa^{2}+24\mu}{\mu^{2}}-3}\quad\Longrightarrow\quad\mu<2+\kappa.$ (4.9)
From it, the perturbation parameter $\mu$ can be written in terms of a chosen value for $n_{\rm max}$ as
$\begin{array}[]{ll}n_{\rm max}=1\!\!:&\quad\displaystyle\frac{2+3\kappa}{7}\leq\mu<2+\kappa.\\[10.0pt] n_{\rm max}>1\!\!:&\quad\displaystyle\frac{2+\kappa(2n_{\rm max}+1)}{1+3n_{\rm max}(n_{\rm max}+1)}\leq\mu<\frac{2+\kappa(2n_{\rm max}-1)}{1+3n_{\rm max}(n_{\rm max}-1)}.\end{array}$
In this way, the expressions (4.8) and (4.9) characterize a cubic hyperbolic perturbation of the spherical oscillator that leads to a positive finite discrete spectrum with ${n}\in\{1,\dots,n_{\rm max}\}$. Observe that the unbounded discrete spectrum (3.11) is recovered under the limit $\mu\to 0$ (so $n_{\rm max}\to\infty$). The flat limit $\kappa\to 0$ is also well-defined, reproducing the hyperbolic perturbation of the Euclidean oscillator determined by (4.6) and (4.7). Figure 5 shows the discrete spectra for several values of $n_{\rm max}$ (so for $\mu$), which can be compared with figure 1 ($\mu=0$) and figure 4 ($\kappa=0$).
Likewise, both types of perturbations can be applied to the oscillator on the hyperbolic space $\mathbf{H}^{2}$ with negative curvature $\kappa=-|\kappa|<0$ (see (3.13)–(3.15)), namely
[…]
A similar analysis as above would be necessary for each type of perturbation. However, the results are quite cumbersome, as there are several possibilities for $n_{\rm max}$ depending on $|\kappa|$ and $\mu$. For brevity in the exposition, these explicit computations are omitted.
[…]
5 The quartic quantum Zernike Hamiltonian
We proceed similarly to the two previous sections, first giving the quantum symmetries and the possible spectra for the quartic Hamiltonian ^$\hat{\mathcal{H}}_{4}$ in (2.3), depending on four arbitrary parameters $\gamma_{1}$, $\gamma_{2}$, $\gamma_{3}$ and $\gamma_{4}$, and imposing that the quantum quadratic Hamiltonian ^$\hat{\mathcal{H}}_{2}$ in (2.3) is recovered for the vanishing of $\gamma_{3}$ and $\gamma_{4}$.
The two quantum symmetries (2.6) of ^$\hat{\mathcal{H}}_{4}$, now of fourth-order in the momenta, read
$\begin{split}{\hat{I}}_{4}&={\hat{p}}_{2}^{2}+\gamma_{1}{\hat{q}}_{2}{\hat{p}}_{2}+\gamma_{2}\bigl{(}({\hat{q}}_{1}^{2}+{\hat{q}}_{2}^{2}){\hat{p}}_{2}^{2}-\hat{\mathcal{C}}^{2}\bigr{)}\\[2.0pt] &\qquad+\gamma_{3}\bigl{(}{\hat{q}}_{2}^{3}({\hat{p}}_{2}^{3}-{\hat{p}}_{1}^{2}{\hat{p}}_{2})+({\hat{q}}_{1}^{3}+3{\hat{q}}_{1}{\hat{q}}_{2}^{2}){\hat{p}}_{1}{\hat{p}}_{2}^{2}-3{\rm i}{\hat{q}}_{2}^{2}{\hat{p}}_{2}^{2}-3{\rm i}{\hat{q}}_{1}{\hat{q}}_{2}{\hat{p}}_{1}{\hat{p}}_{2}-{\hat{q}}_{2}{\hat{p}}_{2}\bigr{)}\\[2.0pt] &\qquad+\gamma_{4}\bigl{(}({\hat{q}}_{2}^{4}-{\hat{q}}_{1}^{4})({\hat{p}}_{2}^{4}-{\hat{p}}_{1}^{2}{\hat{p}}_{2}^{2})+4({\hat{q}}_{1}^{3}{\hat{q}}_{2}+{\hat{q}}_{1}{\hat{q}}_{2}^{3}){\hat{p}}_{1}{\hat{p}}_{2}^{3}-6{\rm i}({\hat{q}}_{2}^{3}+{\hat{q}}_{1}^{2}{\hat{q}}_{2}){\hat{p}}_{2}^{3}\\[2.0pt] &\qquad\qquad\qquad-6{\rm i}({\hat{q}}_{1}^{3}+{\hat{q}}_{1}{\hat{q}}_{2}^{2}){\hat{p}}_{1}{\hat{p}}_{2}^{2}-4({\hat{q}}_{1}^{2}+{\hat{q}}_{2}^{2}){\hat{p}}_{2}^{2}+4\hat{\mathcal{C}}^{2}\bigr{)},\\[4.0pt] {\hat{I}}^{\prime}_{4}&={\hat{p}}_{1}^{2}+\gamma_{1}{\hat{q}}_{1}{\hat{p}}_{1}+\gamma_{2}\bigl{(}{\hat{q}}_{1}^{2}+{\hat{q}}_{2}^{2}\bigr{)}{\hat{p}}_{1}^{2}\\[2.0pt] &\qquad+\gamma_{3}\bigl{(}{\hat{q}}_{1}^{3}({\hat{p}}_{1}^{3}-{\hat{p}}_{1}{\hat{p}}_{2}^{2})+({\hat{q}}_{2}^{3}+3{\hat{q}}_{1}^{2}{\hat{q}}_{2}){\hat{p}}_{1}^{2}{\hat{p}}_{2}-3{\rm i}{\hat{q}}_{1}^{2}{\hat{p}}_{1}^{2}-3{\rm i}{\hat{q}}_{1}{\hat{q}}_{2}{\hat{p}}_{1}{\hat{p}}_{2}-{\hat{q}}_{1}{\hat{p}}_{1}\bigr{)}\\[2.0pt] &\qquad+\gamma_{4}\bigl{(}({\hat{q}}_{1}^{4}-{\hat{q}}_{2}^{4})({\hat{p}}_{1}^{4}-{\hat{p}}_{1}^{2}{\hat{p}}_{2}^{2})+4({\hat{q}}_{1}{\hat{q}}_{2}^{3}+{\hat{q}}_{1}^{3}{\hat{q}}_{2}){\hat{p}}_{1}^{3}{\hat{p}}_{2}-6{\rm i}({\hat{q}}_{1}^{3}+{\hat{q}}_{1}{\hat{q}}_{2}^{2}){\hat{p}}_{1}^{3}\\[2.0pt] &\qquad\qquad\qquad-6{\rm i}({\hat{q}}_{2}^{3}+{\hat{q}}_{1}^{2}{\hat{q}}_{2}){\hat{p}}_{1}^{2}{\hat{p}}_{2}-4({\hat{q}}_{1}^{2}+{\hat{q}}_{2}^{2}){\hat{p}}_{1}^{2}\bigr{)},\end{split}$
and verify the algebraic dependence relation
^$\hat{\mathcal{H}}_{4}={\hat{I}}_{4}+{\hat{I}}_{4}^{\prime}-4\gamma_{4}\hat{\mathcal{C}}^{2}+\gamma_{4}\hat{\mathcal{C}}^{4}.$
Note that, in this case, the classical functions $I_{4}$ and $I_{4}^{\prime}$ are related to $\mathcal{I}_{4}$ and $\mathcal{I}_{4}^{\prime}$ obtained in [3] through $I_{4}=\mathcal{I}_{4}-\gamma_{2}\mathcal{C}^{2}+4\gamma_{4}\mathcal{C}^{2}$ and $I_{4}^{\prime}=\mathcal{I}_{4}^{\prime}$.
The operators $(\hat{\mathcal{K}}_{1},\hat{\mathcal{K}}_{2},\hat{\mathcal{K}}_{3})$ (2.8) fulfil the polynomial commutation relations given by
$\begin{split}\bigl{[}\hat{\mathcal{K}}_{1},\hat{\mathcal{K}}_{2}\bigr{]}&=\hat{\mathcal{K}}_{3},\qquad\bigl{[}\hat{\mathcal{K}}_{1},\hat{\mathcal{K}}_{3}\bigr{]}=4\hat{\mathcal{K}}_{2}-2(\gamma_{2}-4\gamma_{4})\hat{\mathcal{K}}_{1}^{2},\\[2.0pt] \bigl{[}\hat{\mathcal{K}}_{2},\hat{\mathcal{K}}_{3}\bigr{]}&=\bigl{(}\gamma_{1}^{2}+2{\rm i}\gamma_{1}\gamma_{2}-4\gamma_{1}\gamma_{3}-8{\rm i}\gamma_{1}\gamma_{4}+2(\gamma_{2}+3{\rm i}\gamma_{3}-8\gamma_{4})\hat{\mathcal{H}}_{4}\bigr{)}\hat{\mathcal{K}}_{1}\\[2.0pt] &\qquad+4(\gamma_{2}-4\gamma_{4})\hat{\mathcal{K}}_{1}\hat{\mathcal{K}}_{2}-2(\gamma_{2}-4\gamma_{4})\hat{\mathcal{K}}_{3}\\[2.0pt] &\qquad+4\bigl{(}\gamma_{3}^{2}-\gamma_{1}\gamma_{3}+{\rm i}\gamma_{2}\gamma_{3}-8\gamma_{4}^{2}-3{\rm i}\gamma_{1}\gamma_{4}+2{\rm i}\gamma_{3}\gamma_{4}-\gamma_{4}\hat{\mathcal{H}}_{4}\bigr{)}\hat{\mathcal{K}}_{1}^{3}\\[2.0pt] &\qquad+\bigl{(}3\gamma_{3}^{2}+16\gamma_{4}^{2}-6\gamma_{2}\gamma_{4}+6{\rm i}\gamma_{3}\gamma_{4}\bigr{)}\hat{\mathcal{K}}_{1}^{5}+4\gamma_{4}^{2}\hat{\mathcal{K}}_{1}^{7}.\end{split}$
The number and ladder operators $(\hat{\mathcal{K}},\hat{\mathcal{K}}_{+},\hat{\mathcal{K}}_{-})$ are now defined by
$\begin{split}\hat{\mathcal{K}}&:=\frac{1}{2}\hat{\mathcal{K}}_{1},\\ \hat{\mathcal{K}}_{+}&:=\hat{\mathcal{K}}_{2}+\frac{1}{2}\hat{\mathcal{K}}_{3}-\biggl{(}\frac{1}{2}\gamma_{2}-2\gamma_{4}\biggr{)}\hat{\mathcal{K}}_{1}^{2},\\ \hat{\mathcal{K}}_{-}&:=\hat{\mathcal{K}}_{2}-\frac{1}{2}\hat{\mathcal{K}}_{3}-\biggl{(}\frac{1}{2}\gamma_{2}-2\gamma_{4}\biggr{)}\hat{\mathcal{K}}_{1}^{2},\end{split}$ (5.1)
and close on the commutation relations in (2.9) with
$\begin{split}\bigl{[}\hat{\mathcal{K}}_{-},\hat{\mathcal{K}}_{+}\bigr{]}&=2\bigl{(}\gamma_{1}^{2}+2{\rm i}\gamma_{1}\gamma_{2}-4\gamma_{1}\gamma_{3}-8{\rm i}\gamma_{1}\gamma_{4}+2(\gamma_{2}+3{\rm i}\gamma_{3}-8\gamma_{4})\hat{\mathcal{H}}_{4}\bigr{)}\hat{\mathcal{K}}\\[2.0pt] &\qquad+16\bigl{(}\gamma_{2}^{2}+2\gamma_{3}^{2}-2\gamma_{1}\gamma_{3}-6{\rm i}\gamma_{1}\gamma_{4}+2{\rm i}\gamma_{2}\gamma_{3}-8\gamma_{2}\gamma_{4}+4{\rm i}\gamma_{3}\gamma_{4}-2\gamma_{4}\hat{\mathcal{H}}_{4}\bigr{)}\hat{\mathcal{K}}^{3}\\[2.0pt] &\qquad+32\bigl{(}3\gamma_{3}^{2}+16\gamma_{4}^{2}-6\gamma_{2}\gamma_{4}+6{\rm i}\gamma_{3}\gamma_{4}\bigr{)}\hat{\mathcal{K}}^{5}+512\gamma_{4}^{2}\hat{\mathcal{K}}^{7},\end{split}$
hence determining a seventh-order polynomial symmetry algebra of Higgs-type $\mathfrak{sl}^{(7)}(2,\mathbb{R})$. The structure function $\Phi$ in (2.9) turns out to be
$\begin{split}&\hat{\mathcal{K}}_{+}\hat{\mathcal{K}}_{-}=\Phi(\hat{\mathcal{H}}_{4},\hat{\mathcal{K}})\\ &=\frac{1}{4}\bigl{(}4\gamma_{2}-2{\rm i}\gamma_{1}+8{\rm i}\gamma_{3}-16\gamma_{4}+\hat{\mathcal{H}}_{4}\bigr{)}\hat{\mathcal{H}}_{4}\\ &\qquad-\bigl{(}\gamma_{1}^{2}+2{\rm i}\gamma_{1}\gamma_{2}-4\gamma_{1}\gamma_{3}-8{\rm i}\gamma_{1}\gamma_{4}+2(\gamma_{2}+3{\rm i}\gamma_{3}-8\gamma_{4})\hat{\mathcal{H}}_{4}\bigr{)}\hat{\mathcal{K}}\\[2.0pt] &\qquad+\bigl{(}\gamma_{1}^{2}+4\gamma_{2}^{2}+2{\rm i}\gamma_{1}\gamma_{2}-12\gamma_{1}\gamma_{3}-32{\rm i}\gamma_{1}\gamma_{4}+8{\rm i}\gamma_{2}\gamma_{3}-16\gamma_{2}\gamma_{4}\\[2.0pt] &\qquad\qquad+2(\gamma_{2}+3{\rm i}\gamma_{3}-12\gamma_{4})\hat{\mathcal{H}}_{4}\bigr{)}\hat{\mathcal{K}}^{2}\\[2.0pt] &\qquad-8\bigl{(}\gamma_{2}^{2}+2\gamma_{3}^{2}-2\gamma_{1}\gamma_{3}-6{\rm i}\gamma_{1}\gamma_{4}+2{\rm i}\gamma_{2}\gamma_{3}-8\gamma_{2}\gamma_{4}+4{\rm i}\gamma_{3}\gamma_{4}-2\gamma_{4}\hat{\mathcal{H}}_{4}\bigr{)}\hat{\mathcal{K}}^{3}\\[2.0pt] &\qquad+4\bigl{(}\gamma_{2}^{2}+12\gamma_{3}^{2}+16\gamma_{4}^{2}-2\gamma_{1}\gamma_{3}-6{\rm i}\gamma_{1}\gamma_{4}+2{\rm i}\gamma_{2}\gamma_{3}-28\gamma_{2}\gamma_{4}+24{\rm i}\gamma_{3}\gamma_{4}-2\gamma_{4}\hat{\mathcal{H}}_{4}\bigr{)}\hat{\mathcal{K}}^{4}\\[2.0pt] &\qquad-16\bigl{(}3\gamma_{3}^{2}+16\gamma_{4}^{2}-6\gamma_{2}\gamma_{4}+6{\rm i}\gamma_{3}\gamma_{4}\bigr{)}\hat{\mathcal{K}}^{5}\\[2.0pt] &\qquad+16\bigl{(}\gamma_{3}^{2}+24\gamma_{4}^{2}-2\gamma_{2}\gamma_{4}+2{\rm i}\gamma_{3}\gamma_{4}\bigr{)}\hat{\mathcal{K}}^{6}-256\gamma_{4}^{2}\hat{\mathcal{K}}^{7}+64\gamma_{4}^{2}\hat{\mathcal{K}}^{8},\end{split}$
which further factorizes as
$\begin{split}\Phi(\hat{\mathcal{H}}_{4},\hat{\mathcal{K}})&=\Phi_{1}(\hat{\mathcal{H}}_{4},\hat{\mathcal{K}})\Phi_{2}(\hat{\mathcal{H}}_{4},\hat{\mathcal{K}}),\qquad\bigl{[}\Phi_{1}(\hat{\mathcal{H}}_{4},\hat{\mathcal{K}}),\Phi_{2}(\hat{\mathcal{H}}_{4},\hat{\mathcal{K}})\bigr{]}=0,\\ \Phi_{1}(\hat{\mathcal{H}}_{4},\hat{\mathcal{K}})&=\frac{1}{4}\bigl{(}\hat{\mathcal{H}}_{4}-2{\rm i}\gamma_{1}\hat{\mathcal{K}}+4\gamma_{2}\hat{\mathcal{K}}^{2}+8{\rm i}\gamma_{3}\hat{\mathcal{K}}^{3}-16\gamma_{4}\hat{\mathcal{K}}^{4}\bigr{)},\\[2.0pt] \Phi_{2}(\hat{\mathcal{H}}_{4},\hat{\mathcal{K}})&=\hat{\mathcal{H}}_{4}-2({\rm i}\gamma_{1}-2\gamma_{2}-4{\rm i}\gamma_{3}+8\gamma_{4}){\rm Id}+2({\rm i}\gamma_{1}-4\gamma_{2}-12{\rm i}\gamma_{3}+32\gamma_{4})\hat{\mathcal{K}}\\[2.0pt] &\qquad+4(\gamma_{2}+6{\rm i}\gamma_{3}-24\gamma_{4})\hat{\mathcal{K}}^{2}-8({\rm i}\gamma_{3}-8\gamma_{4})\hat{\mathcal{K}}^{3}-16\gamma_{4}\hat{\mathcal{K}}^{4}\\[2.0pt] &=\hat{\mathcal{H}}_{4}+2{\rm i}\gamma_{1}\bigl{(}\hat{\mathcal{K}}-{\rm Id}\bigr{)}+4\gamma_{2}\bigl{(}\hat{\mathcal{K}}-{\rm Id}\bigr{)}^{2}-8{\rm i}\gamma_{3}\bigl{(}\hat{\mathcal{K}}-{\rm Id}\bigr{)}^{3}-16\gamma_{4}\bigl{(}\hat{\mathcal{K}}-{\rm Id}\bigr{)}^{4}.\end{split}$ (5.2)
The change of operators (2.11) leads to the deformed oscillator algebra (2.12).
5.1 Spectrum of the quartic system
We introduce the finite-dimensional representation of (2.12) into the expressions (5.2), obtaining the structure function $\Phi=\Phi_{1}\Phi_{2}$ (2.13) as
$\begin{split}\Phi_{1}({B},E,u)&=\frac{1}{4}\bigl{(}E-2{\rm i}\gamma_{1}({B}+u)+4\gamma_{2}({B}+u)^{2}+8{\rm i}\gamma_{3}({B}+u)^{3}-16\gamma_{4}({B}+u)^{4}\bigr{)},\\[2.0pt] \Phi_{2}({B},E,u)&=E+2{\rm i}\gamma_{1}({B}+u-1)+4\gamma_{2}({B}+u-1)^{2}-8{\rm i}\gamma_{3}({B}+u-1)^{3}-16\gamma_{4}({B}+u-1)^{4}.\end{split}$ (5.3)
The spectrum $E$ of ^$\hat{\mathcal{H}}_{4}$ (2.4) is derived by imposing the two conditions (2.14) and is established as follows.
Proposition 3
(i) The set of equations (2.14) obtained from (5.3) leads to twelve types of solutions for the representation dependent constant $u=u({n})$ and the spectrum $E=E({n})$ of the Hamiltonian ^$\hat{\mathcal{H}}_{4}$ (2.4) depending on the parameter ${n}\in\{1,2,\dots\}$. They provide the structure function $\Phi({B},E({n}),u({n}))\equiv\Phi({B},{n})$ for ${B}=\{1,\dots,{n}\}$ through (5.3).
(ii) Among all these solutions, there are only two types such that the limits $\gamma_{4}\to 0$ and $\gamma_{3}\to 0$ are simultaneously well-defined for any value of $\gamma_{1}$ and $\gamma_{2}$, namely
[…]
The ten excluded solutions exhibit divergencies, similarly to the solution (4.2) for the cubic Hamiltonian ^$\hat{\mathcal{H}}_{3}$, although the resulting expressions are obviously more complicated in this case (they involve both parameters $\gamma_{3}$ and $\gamma_{4}$ within the denominators). In this sense, the discarded solutions in propositions 2 and 3 can be regarded as ‘spurious’.
It is worth remarking that the results given in proposition 3 are well-defined for any value of the four parameters $\gamma_{k}$ $(k=1,\dots,4)$. As expected, they lead exactly to the results presented in proposition 2 for $\gamma_{4}=0$ and, then, to those characterized by proposition 1 when $\gamma_{3}=0$. In addition, it is possible to set $\gamma_{3}=0$ in the solutions (5.4), keeping the parameter $\gamma_{4}$, which determines a superintegrable quartic perturbation of the initial Hamiltonian ^$\hat{\mathcal{H}}_{2}$.
We observe that the expressions (5.4) take real values whenever the odd coefficients $\gamma_{1}$ and $\gamma_{3}$ are pure imaginary numbers, while the even ones $\gamma_{2}$ and $\gamma_{4}$ take real values. This difference between odd/even parameters has already been studied in the classical picture [3]. Therefore, by taking into account (4.3), let us denote
$\gamma_{1}=-\rm i\beta,\qquad\gamma_{2}=\alpha,\qquad\gamma_{3}={\rm i}\mu,\qquad\gamma_{4}=-\nu,\qquad\{\alpha,\beta,\mu,\nu\}\in\mathbb{R},$
obtaining two types of spectra for ^$\hat{\mathcal{H}}_{4}$ given by
$\begin{split}E_{\rm I}&=-\bigl{(}\beta{n}+\alpha{n}^{2}+\mu{n}^{3}+\nu{n}^{4}\bigr{)},\\[2.0pt] E_{\rm II}&=\beta({n}+2)-\alpha({n}+2)^{2}+\mu({n}+2)^{3}-\nu({n}+2)^{4},\end{split}$
to be compared with (4.4). In addition, if we set $\alpha=-1$ and $\beta=-2$ and interpret ${n}$ as the principal quantum number $\mathfrak{n}$, we find two possible superintegrable perturbations of the spectrum (3.8) of the Zernike Hamiltonian ^$\hat{\mathcal{H}}_{\rm Zk}$ in (1.1) depending on two real parameters $\mu$ and $\nu$, and given by
$E_{\rm I}=\mathfrak{n}(\mathfrak{n}+2)-\mu\mathfrak{n}^{3}-\nu\mathfrak{n}^{4},\qquad E_{\rm II}=\mathfrak{n}(\mathfrak{n}+2)+\mu(\mathfrak{n}+2)^{3}-\nu(\mathfrak{n}+2)^{4}.$
Finally, we remark that the results of proposition 3 can be applied to the isotropic oscillator on $\mathbf{E}^{2}$, $\mathbf{S}^{2}$ and $\mathbf{H}^{2}$, described in section 3.2, thus obtaining for them different combinations of quartic/cubic perturbations of spherical/hyperbolic type following a similar analysis as in section 4.2.
6 Higher-order quantum Zernike Hamiltonians: Conjectures
Although the quantum Hamiltonian ^$\hat{\mathcal{H}}_{N}$ has been analyzed in detail for the values $N<5$, a closed explicit expression for the higher-order quantum symmetries ${\hat{I}}_{N}$ and ${\hat{I}}^{\prime}_{N}$ for arbitrary values of $N$ has not been found, due to the exponentially increasing computational difficulties. However, based on the results for the values analyzed, the form of the structure function $\Phi(\hat{\mathcal{H}}_{N},\hat{\mathcal{K}})$ in (2.9) along with its factorizing terms (2.10) can be inferred, allowing us to deduce the two types I and II of solutions for the spectrum that generalize those in (5.4) and preserve both the superintegrability and the properties concerning the limits $\gamma_{k}\rightarrow 0$ for $k\geq 3$. This extrapolation is presented in form of conjectures, and to illustrate their validity, we summarize the case for $N=5$, which has been explicitly solved, and indicate some of the computational obstructions to be expected for large values of $N$.
Conjecture 1
(i) For any $N$ and any value of the coefficients $\gamma_{k}$ ($1\leq k\leq N)$, the quantum Hamiltonian ^$\hat{\mathcal{H}}_{N}$ (2.3) admits three quantum symmetries corresponding to the quantum angular momentum operator $\hat{\mathcal{C}}$ in (2.5) and two operators ${\hat{I}}_{N}$ and ${\hat{I}}^{\prime}_{N}$ (2.6) which are quadratic in the momenta for $N=1,2$ and of $N^{{\rm th}}$-order for $N>2$.
(ii) The sets {^$\{\hat{\mathcal{H}}_{N},\hat{\mathcal{C}},{\hat{I}}_{N}\}$ and {^$\{\hat{\mathcal{H}}_{N},\hat{\mathcal{C}},{\hat{I}}^{\prime}_{N}\}$ are formed by three algebraically independent operators, hence ^$\hat{\mathcal{H}}_{N}$ determines a superintegrable system.
(iii) The set $\{\hat{\mathcal{C}},{\hat{I}}_{N},{\hat{I}}^{\prime}_{N}\}$ leads to the quantum symmetries $(\hat{\mathcal{K}}_{1},\hat{\mathcal{K}}_{2},\hat{\mathcal{K}}_{3})$ in (2.8) and, in consequence, it is possible to define the number and ladder operators $(\hat{\mathcal{K}},\hat{\mathcal{K}}_{+},\hat{\mathcal{K}}_{-})$ verifying the commutation relations (2.9) with the structure function $\Phi=\Phi_{1}\Phi_{2}$ (2.10), whose factorizing terms are given by
$\begin{split}&\Phi_{1}\bigl{(}\hat{\mathcal{H}}_{N},\hat{\mathcal{K}}\bigr{)}=\frac{1}{4}\left(\hat{\mathcal{H}}_{N}-\sum_{k=1}^{N}(2{\rm i})^{k}\gamma_{k}\hat{\mathcal{K}}^{k}\right),\\ &\Phi_{2}\bigl{(}\hat{\mathcal{H}}_{N},\hat{\mathcal{K}}\bigr{)}=\hat{\mathcal{H}}_{N}-\sum_{k=1}^{N}(-2{\rm i})^{k}\gamma_{k}\bigl{(}\hat{\mathcal{K}}-{\rm Id}\bigr{)}^{k}.\end{split}$ (6.1)
Observe that $[\Phi_{1},\Phi_{2}]=0$ and that the operators $(\hat{\mathcal{K}},\hat{\mathcal{K}}_{+},\hat{\mathcal{K}}_{-})$ should span a $(2N-1)^{{\rm th}}$-order polynomial symmetry algebra of Higgs-type $\mathfrak{sl}^{(2N-1)}(2,\mathbb{R})$ determined by the commutator $[\hat{\mathcal{K}}_{-},\hat{\mathcal{K}}_{+}]$.
Provided that the conjecture 1 is satisfied, we can further deduce the factorizing terms in (2.10) (hence the structure function $\Phi$) and the solutions consistent with the generalization of the quantum Zernike Hamiltonian.
Conjecture 2
(i) The structure function $\Phi({B},E,u)=\Phi_{1}\Phi_{2}$ (2.13) is obtained from (6.1) by introducing the finite-dimensional representation of (2.12), namely
$\begin{split}&\Phi_{1}({B},E,u)=\frac{1}{4}\left(E-\sum_{k=1}^{N}(2{\rm i})^{k}\gamma_{k}(B+u)^{k}\right),\\ &\Phi_{2}({B},E,u)=E-\sum_{k=1}^{N}(-2{\rm i})^{k}\gamma_{k}(B+u-1)^{k}.\end{split}$ (6.2)
(ii) The set of equations (2.14) coming from (6.2) gives rise to two solutions for the representation dependent constant $u=u({n})$ and the spectrum $E=E({n})$ of the Hamiltonian ^$\hat{\mathcal{H}}_{N}$ (2.4) depending on the parameter ${n}\in\{1,2,\dots\}$, that hold for any values of the coefficients $\gamma_{k}$. From these solutions the structure function $\Phi({B},E({n}),u({n}))\equiv\Phi({B},{n})=\Phi_{1}({B},{n})\Phi_{2}({B},{n})$ for ${B}=\{1,\dots,{n}\}$ is obtained using (6.2). The resulting expressions read as
[…]
Some observations concerning conjecture 2 seem pertinent. First of all, it can be easily checked that both types of solutions satisfy the equations (2.14), as
$\Phi_{1}(0,E_{\rm I},u_{\rm I})=\Phi_{2}(n+1,E_{\rm I},u_{\rm I})=0,\qquad\Phi_{1}(n+1,E_{\rm II},u_{\rm II})=\Phi_{2}(0,E_{\rm II},u_{\rm II})=0.$
Second, to ensure that both spectra $E_{\rm I}$ and $E_{\rm II}$ are real, it is necessary to require that $\gamma_{k}$ is purely imaginary for odd values of $k$, while it must be real for even values of $k$. Third, the above results are well-defined for any value of all the coefficients $\gamma_{k}$. In particular, we can take a single or more coefficients different from zero. For instance, the simplest choice is to set only $\gamma_{1}\neq 0$, leading to a linear spectrum in $n$, that corresponds to the first system with a discrete spectrum; in fact, it has been interpreted as the isotropic Euclidean oscillator in section 3.2 (when $\gamma_{k}=0$ for all $k$, we recover the trivial geodesic motion on the Euclidean plane). Finally, from this point of view, the generalized quantum Zernike Hamiltonian ^$\hat{\mathcal{H}}_{N}$ in (2.3) can be regarded as the superposition of $N$ possible potentials, each of them determined by one $\gamma_{k}\neq 0$, leading to a term of $k^{{\rm th}}$-order in the momenta in the quantum symmetries ${\hat{I}}_{N}$ and ${\hat{I}}^{\prime}_{N}$, as well as a term $n^{k}$ (or $(n+2)^{k}$) in the spectrum.
We now outline the main results for the case $N=5$, which are in full agreement with both conjectures. The quantum symmetry ${\hat{I}}^{\prime}_{5}$ in (2.6) of ^$\hat{\mathcal{H}}_{5}$ is explicitly obtained, after some cumbersome computations, as
$\begin{split}{\hat{I}}^{\prime}_{5}&={\hat{p}}_{1}^{2}+\gamma_{1}{\hat{q}}_{1}{\hat{p}}_{1}+\gamma_{2}\bigl{(}{\hat{q}}_{1}^{2}+{\hat{q}}_{2}^{2}\bigr{)}{\hat{p}}_{1}^{2}\\[2.0pt] &\qquad+\gamma_{3}\bigl{(}{\hat{q}}_{1}^{3}({\hat{p}}_{1}^{3}-{\hat{p}}_{1}{\hat{p}}_{2}^{2})+({\hat{q}}_{2}^{3}+3{\hat{q}}_{1}^{2}{\hat{q}}_{2}){\hat{p}}_{1}^{2}{\hat{p}}_{2}-3{\rm i}{\hat{q}}_{1}^{2}{\hat{p}}_{1}^{2}-3{\rm i}{\hat{q}}_{1}{\hat{q}}_{2}{\hat{p}}_{1}{\hat{p}}_{2}-{\hat{q}}_{1}{\hat{p}}_{1}\bigr{)}\\[2.0pt] &\qquad+\gamma_{4}\bigl{(}({\hat{q}}_{1}^{4}-{\hat{q}}_{2}^{4})({\hat{p}}_{1}^{4}-{\hat{p}}_{1}^{2}{\hat{p}}_{2}^{2})+4({\hat{q}}_{1}{\hat{q}}_{2}^{3}+{\hat{q}}_{1}^{3}{\hat{q}}_{2}){\hat{p}}_{1}^{3}{\hat{p}}_{2}-6{\rm i}({\hat{q}}_{1}^{3}+{\hat{q}}_{1}{\hat{q}}_{2}^{2}){\hat{p}}_{1}^{3}\\[2.0pt] &\qquad\qquad\quad-6{\rm i}({\hat{q}}_{2}^{3}+{\hat{q}}_{1}^{2}{\hat{q}}_{2}){\hat{p}}_{1}^{2}{\hat{p}}_{2}-4({\hat{q}}_{1}^{2}+{\hat{q}}_{2}^{2}){\hat{p}}_{1}^{2}\bigr{)}\\[2.0pt] &\qquad+\gamma_{5}\bigl{(}{\hat{q}}_{1}^{5}({\hat{p}}_{1}^{5}+{\hat{p}}_{1}{\hat{p}}_{2}^{4})-({\hat{q}}_{2}^{5}-5{\hat{q}}_{1}^{4}{\hat{q}}_{2})({\hat{p}}_{1}^{4}{\hat{p}}_{2}-{\hat{p}}_{1}^{2}{\hat{p}}_{2}^{3})-({\hat{q}}_{1}^{5}-10{\hat{q}}_{1}^{3}{\hat{q}}_{2}^{2}-5{\hat{q}}_{1}{\hat{q}}_{2}^{4}){\hat{p}}_{1}^{3}{\hat{p}}_{2}^{2}\\[2.0pt] &\qquad\qquad\quad-10{\rm i}{\hat{q}}_{1}^{4}({\hat{p}}_{1}^{4}-{\hat{p}}_{1}^{2}{\hat{p}}_{2}^{2})-10{\rm i}({\hat{q}}_{1}{\hat{q}}_{2}^{3}+4{\hat{q}}_{1}^{3}{\hat{q}}_{2}){\hat{p}}_{1}^{3}{\hat{p}}_{2}-10{\rm i}({\hat{q}}_{2}^{4}+3{\hat{q}}_{1}^{2}{\hat{q}}_{2}^{2}){\hat{p}}_{1}^{2}{\hat{p}}_{2}^{2}\\[2.0pt] &\qquad\qquad\quad+10{\rm i}{\hat{q}}_{1}^{3}{\hat{q}}_{2}{\hat{p}}_{1}{\hat{p}}_{2}^{3}-25{\hat{q}}_{1}^{3}{\hat{p}}_{1}^{3}-10({\hat{q}}_{2}^{3}+6{\hat{q}}_{1}^{2}{\hat{q}}_{2}){\hat{p}}_{1}^{2}{\hat{p}}_{2}+5(2{\hat{q}}_{1}^{3}-3{\hat{q}}_{1}{\hat{q}}_{2}^{2}){\hat{p}}_{1}{\hat{p}}_{2}^{2}\\[2.0pt] &\qquad\qquad\quad+15{\rm i}{\hat{q}}_{1}^{2}{\hat{p}}_{1}^{2}+15{\rm i}{\hat{q}}_{1}{\hat{q}}_{2}{\hat{p}}_{1}{\hat{p}}_{2}+{\hat{q}}_{1}{\hat{p}}_{1}\bigr{)}.\end{split}$
We omit the detailed expression for ${\hat{I}}_{5}$, as it can be deduced from the algebraic dependence relation
^$\hat{\mathcal{H}}_{5}={\hat{I}}_{5}+{\hat{I}}_{5}^{\prime}-4\gamma_{4}\hat{\mathcal{C}}^{2}+\gamma_{4}\hat{\mathcal{C}}^{4}.$
Next we introduce the operators $(\hat{\mathcal{K}}_{1},\hat{\mathcal{K}}_{2},\hat{\mathcal{K}}_{3})$ in (2.8), from which the number and ladder operators $(\hat{\mathcal{K}},\hat{\mathcal{K}}_{+},\hat{\mathcal{K}}_{-})$ with the same expression as in (5.1) can be defined. These operators close a polynomial symmetry algebra $\mathfrak{sl}^{(9)}(2,\mathbb{R})$ fulfilling the commutation rules (2.9), in such a manner that the factorizing terms in (2.10) are given by
$\begin{split}\Phi_{1}&=\frac{1}{4}\bigl{(}\hat{\mathcal{H}}_{5}-2{\rm i}\gamma_{1}\hat{\mathcal{K}}+4\gamma_{2}\hat{\mathcal{K}}^{2}+8{\rm i}\gamma_{3}\hat{\mathcal{K}}^{3}-16\gamma_{4}\hat{\mathcal{K}}^{4}-32{\rm i}\gamma_{5}\hat{\mathcal{K}}^{5}\bigr{)},\\[2.0pt] \Phi_{2}&=\hat{\mathcal{H}}_{5}+2{\rm i}\gamma_{1}\bigl{(}\hat{\mathcal{K}}-{\rm Id}\bigr{)}+4\gamma_{2}\bigl{(}\hat{\mathcal{K}}-{\rm Id}\bigr{)}^{2}-8{\rm i}\gamma_{3}\bigl{(}\hat{\mathcal{K}}-{\rm Id}\bigr{)}^{3}-16\gamma_{4}\bigl{(}\hat{\mathcal{K}}-{\rm Id}\bigr{)}^{4}+32{\rm i}\gamma_{5}\bigl{(}\hat{\mathcal{K}}-{\rm Id}\bigr{)}^{5}.\end{split}$
From these we directly get the relations (6.2) for $N=5$, and thus the solutions to equations (2.14). The two admissible possibilities for the spectrum that emerge imposing the constraint that they hold for any value of the coefficients $\gamma_{k}$ ($1\leq k\leq 5$) are
Type $\begin{array}[]{lll}\mbox{\rm Type I:}&\displaystyle u_{\rm I}=-\frac{{n}}{2},&\quad\displaystyle E_{\rm I}=-\bigl{(}{\rm i}\gamma_{1}{n}+\gamma_{2}{n}^{2}-{\rm i}\gamma_{3}{n}^{3}-\gamma_{4}{n}^{4}+{\rm i}\gamma_{5}n^{5}\bigr{)}.\\[10.0pt] \mbox{\rm Type II:}&\displaystyle u_{\rm II}=-\frac{{n}}{2},&\quad E_{\rm II}={\rm i}\gamma_{1}({n}+2)-\gamma_{2}({n}+2)^{2}-{\rm i}\gamma_{3}({n}+2)^{3}+\gamma_{4}({n}+2)^{4}+{\rm i}\gamma_{5}(n+2)^{5}.\end{array}$
These formulae coincide with those obtained from both conjectures when substituting the value of $N$.
In summary, in order to consistently prove the conjectures 1 and 2, it is first necessary to find the generic expression of the quantum symmetries ${\hat{I}}_{N}$ and ${\hat{I}}^{\prime}_{N}$ in (2.6), from which the structure of the polynomial Higgs-type algebra of order $(2N-1)^{{\rm th}}$ is obtained, which constitutes the main computational obstruction. A second step consists in finding the precise structure of the quantum operators $(\hat{\mathcal{K}},\hat{\mathcal{K}}_{+},\hat{\mathcal{K}}_{-})$, starting from $(\hat{\mathcal{K}}_{1},\hat{\mathcal{K}}_{2},\hat{\mathcal{K}}_{3})$ in (2.8), leading to the factorization terms in (6.1). As shown above, this is far from being a trivial task, even with the use of symbolic computer packages.
7 Conclusions and outlook
In this work, a quantum version of the generalized classical Zernike Hamiltonian first studied in [3] has been proposed, showing that one of its main features, namely the superintegrability property, is preserved by quantization. In addition, it has been shown that the spectrum of the quantum Hamiltonian (2.3) can be inferred by purely algebraic means, using the quantum symmetries and the structure of the associated polynomial Higgs-type algebra. Although explicit computations become extremely cumbersome for values $N\geq 6$, a generic prescription has been given to solve the spectrum problem for any given value of $N$. Based on the general shape of the factorizing terms of the structure function and the resulting spectra for the admissible solutions that allow us to interpret the result as a higher-order perturbation of the usual quantum Zernike model, we conjecture the solutions for arbitrary value of $N$, without making explicit use of the quantum symmetries, for which the exact expression is still an open problem. In this context, another unanswered question concerns the number of solutions of equation (2.14), and whether for higher values of $N$ we can find additional solutions that are well-defined and behave properly for the limits $\gamma_{k}\rightarrow 0$ for $k\geq 3$, albeit this seems not probable in view of the structure observed for values $N\leq 5$.
Furthermore, we stress that the interpretation of the Hamiltonian $\mathcal{H}_{2}$ (2.3) as defining, in a unified form, the isotropic oscillator on $\mathbf{E}^{2}$, $\mathbf{S}^{2}$ and $\mathbf{H}^{2}$, have allowed us to consider the generalized Zernike systems as superintegrable perturbations for these oscillators when $N\geq 3$. Moreover, we have also shown that such perturbations can be of spherical or hyperbolic type depending on the value of the corresponding coefficient $\gamma_{k}$. In this sense, these new results extend Bertrand’s theorem to the momentum-dependent systems studied here.
In any case, the main open problem is to solve the corresponding Schrödinger equation (2.4) by obtaining both eigenfunctions and spectra, and then analyze them with respect to the algebraic results. This task is quite arduous, as demonstrated by the solutions for $N=2$, already carried out in [35, 38].
There is a number of interesting problems that emerge from our analysis, and that deserve a more profound investigation. In particular:
  • The Schrödinger equation could be faced for the simplest flat Euclidean system at least for $N=3$ and the resulting wave eigenfunctions and spectrum analyzed according to the spherical or hyperbolic type of the perturbation. This result can provide the path to solve the Schrödinger equation for the perturbations of curved oscillators on $\mathbf{S}^{2}$ and $\mathbf{H}^{2}$.
  • The generalized quantum Zernike Hamiltonians could be implemented in the three Lorentzian spaces of constant curvature, Minkowski, anti-de Sitter and de Sitter spacetimes, by means of graded contractions or analytical continuations [40]; the appropriate coordinates would be the geodesic polar ones shown in (3.9).
  • As higher-order momentum-dependent potentials have already been considered in quantum molecular and nuclear dynamics (see [41, 42, 43, 44, 45] and references therein), it would be natural to search for some application of the generalized Zernike systems in these frameworks.
Other possible open problems arise when considering the underlying $\mathfrak{gl}(2)$-coalgebra symmetry of the Hamiltonian $\mathcal{H}_{N}$ (2.3), for any $N$, which we briefly summarize.
Let us consider the Lie algebra $\mathfrak{gl}(2)=\operatorname{span}\{J_{-},J_{+},J_{3},\Xi\}$ with commutation relations and Casimir operator ${C}$ given by
$[J_{3},J_{+}]=2{\rm i}J_{+},\qquad[J_{3},J_{-}]=-2{\rm i}J_{-},\qquad[J_{-},J_{+}]=4{\rm i}J_{3}+2\,\Xi,\qquad[\Xi\,,\,\cdot\,]=0,$ (7.1)
${C}=\frac{1}{2}(J_{+}J_{-}+J_{-}J_{+})-J_{3}^{2}+{\rm i}J_{3}\,\Xi+\Xi^{2}.$ (7.2)
Thus, $\Xi$ is a trivial central generator such that $\mathfrak{gl}(2)\simeq\mathfrak{sl}(2,\mathbb{R})\oplus\mathbb{R}$. As for any Lie algebra, the coalgebra symmetry $(\mathfrak{gl}(2),\Delta)$ is defined via the primitive or non-deformed coproduct map $\Delta$:
$\Delta:\mathfrak{gl}(2)\to\mathfrak{gl}(2)\otimes\mathfrak{gl}(2),\qquad\Delta(X)=X\otimes 1+1\otimes X,\quad X\in\{J_{-},J_{+},J_{3},\Xi\},$ (7.3)
which is a Lie algebra homomorphism. Recall that the (trivial) counit and antipode can also be defined leading to a non-deformed Hopf algebra structure [46, 47].
A ‘one-particle’ representation $D$ of $\mathfrak{gl}(2)$ (7.1) in terms of the operators $({\hat{q}}_{1},{\hat{p}}_{1})$ (2.1) reads as
$\begin{split}\hat{J}^{(1)}_{-}&=D(J_{-})={\hat{q}}_{1}^{2},\qquad\ \ \hat{J}^{(1)}_{+}=D(J_{+})={\hat{p}}_{1}^{2}+\frac{\lambda_{1}}{{\hat{q}}_{1}^{2}},\\ \hat{J}^{(1)}_{3}&=D(J_{3})={\hat{q}}_{1}{\hat{p}}_{1},\qquad\hat{\Xi}^{(1)}=D(\Xi)=1.\end{split}$
Hence, it depends on a real parameter $\lambda_{1}$ which labels the representation, as the Casimir operator in (7.2) becomes
$\hat{C}^{(1)}=D(C)=\frac{1}{2}\left(\hat{J}^{(1)}_{+}\hat{J}^{(1)}_{-}+\hat{J}^{(1)}_{-}\hat{J}^{(1)}_{+}\right)-\left(\hat{J}^{(1)}_{3}\right)^{2}+{\rm i}\hat{J}_{3}^{(1)}\,\hat{\Xi}^{(1)}+\left(\hat{\Xi}^{(1)}\right)^{2}=\lambda_{1}.$
Then a ‘two-particle’ representation $D^{(2)}=D\otimes D$ of $\mathfrak{gl}(2)$ (7.1) is directly provided by the coproduct (7.3), namely
$\begin{split}\hat{J}^{(2)}_{-}&=D^{(2)}\bigl{(}\Delta(J_{-})\bigr{)}={\hat{q}}_{1}^{2}+{\hat{q}}_{2}^{2},\qquad\ \ \hat{J}^{(2)}_{+}=D^{(2)}\bigl{(}\Delta(J_{+})\bigr{)}={\hat{p}}_{1}^{2}+\frac{\lambda_{1}}{{\hat{q}}_{1}^{2}}+{\hat{p}}_{2}^{2}+\frac{\lambda_{2}}{{\hat{q}}_{2}^{2}},\\ \hat{J}^{(2)}_{3}&=D^{(2)}\bigl{(}\Delta(J_{3})\bigr{)}={\hat{q}}_{1}{\hat{p}}_{1}+{\hat{q}}_{2}{\hat{p}}_{2},\qquad\hat{\Xi}^{(2)}=D^{(2)}\bigl{(}\Delta(\Xi)\bigr{)}=2,\end{split}$ (7.4)
while for the Casimir operator in (7.2), we find that
$\begin{split}\hat{C}^{(2)}&=D^{(2)}\bigl{(}\Delta(C)\bigr{)}=\frac{1}{2}\left(\hat{J}^{(2)}_{+}\hat{J}^{(2)}_{-}+\hat{J}^{(2)}_{-}\hat{J}^{(2)}_{+}\right)-\left(\hat{J}^{(2)}_{3}\right)^{2}+{\rm i}\hat{J}_{3}^{(2)}\,\hat{\Xi}^{(2)}+\left(\hat{\Xi}^{(2)}\right)^{2}\\ &=({\hat{q}_{1}}{\hat{p}_{2}}-{\hat{q}_{2}}{\hat{p}_{1}})^{2}+\lambda_{1}\,\frac{\hat{q}_{2}^{2}}{\hat{q}_{1}^{2}}+\lambda_{2}\,\frac{\hat{q}_{1}^{2}}{\hat{q}_{2}^{2}}+\lambda_{1}+\lambda_{2}+2.\end{split}$ (7.5)
By construction [48, 49], $\hat{C}^{(2)}$ commutes with the operators (7.4), so that any smooth function ${H}$ defined on them determines a two-dimensional quantum integrable Hamiltonian:
$\hat{H}^{(2)}=H\left(\hat{J}_{3}^{(2)},\hat{J}_{+}^{(2)},\hat{J}_{-}^{(2)},\hat{\Xi}^{(2)}\right).$
The coassociative property of the coproduct allows one to extend this result to arbitrary dimension $d$ providing $(2d-3)$ ‘universal’ algebraically independent operators [49] for $\hat{H}^{(d)}$, so that the resulting systems are called quasi-maximally superintegrable since only one additional quantum symmetry is needed to ensure maximal superintegrability. In the classical case, the central generator $\Xi$ vanishes, so that $(\mathfrak{gl}(2),\Delta)$ reduces to a Poisson $\mathfrak{sl}(2,\mathbb{R})$-coalgebra, which has been deeply studied [50, 49]. Recall also that the Racah algebra for $\mathfrak{sl}(2,\mathbb{R})$ has been widely studied in [51, 52, 53].
In our case, let us choose ($d=2$)
$\hat{H}^{(2)}=\hat{J}_{+}^{(2)}+\sum_{k=1}^{N}\gamma_{k}\left(\hat{J}_{3}^{(2)}\right)^{k}=\hat{\mathbf{p}}^{2}+\sum_{k=1}^{N}\gamma_{k}(\hat{\mathbf{q}}\boldsymbol{\cdot}\hat{\mathbf{p}})^{k}+\frac{\lambda_{1}}{\hat{q}_{1}^{2}}+\frac{\lambda_{2}}{\hat{q}_{2}^{2}}.$ (7.6)
Therefore, if we set $\lambda_{1}=\lambda_{2}=0$ we recover the generalized quantum Zernike Hamiltonian ^$\hat{\mathcal{H}}_{N}$ (2.3), while the operator $\hat{C}^{(2)}$ in (7.5) gives the square of the quantum angular momentum operator (2.5). Hence, for arbitrary coefficients $\lambda_{1}$ and $\lambda_{2}$, the Hamiltonian $\hat{H}^{(2)}$ is formed by the superposition of ^$\hat{\mathcal{H}}_{N}$ with two Rosochatius–Winternitz potentials. In the classical case with $N=2$ and within the interpretation of ^$\hat{\mathcal{H}}_{2}$ as the Hamiltonian of a curved oscillator, $\hat{H}^{(2)}$ is just the curved Smorodinsky–Winternitz system which is known to be superintegrable [50]; the presence of a positive $\lambda_{i}$-potential corresponds to introduce a ‘centrifugal’ barrier on $\mathbf{E}^{2}$ and $\mathbf{H}^{2}$, but a noncentral oscillator on $\mathbf{S}^{2}$ [36]. From this perspective, it is rather natural to look for an additional quantum symmetry for $\hat{H}^{(2)}$ (7.6) and to deduce and analyze the resulting spectrum by applying an algebraic approach.
Finally, it would also be possible to consider a quantum deformation of $\mathfrak{gl}(2)$, inherited from a quantum $\mathfrak{sl}(2,\mathbb{R})$ algebra, with quantum deformation parameter $q=\exp z$. In our case, it would be appropriate to make use of the so called non-standard deformation of $\mathfrak{sl}(2,\mathbb{R})$ instead of the usual Drinfel’d–Jimbo one. The reason is that for the former the generator $J_{-}$ remains primitive (with trivial coproduct) and, under the representation (7.4), there appear factors of the type $\exp(z({\hat{q}}_{1}^{2}+{\hat{q}}_{2}^{2}))$ explicitly in the deformed analogue of $\hat{H}^{(2)}$ (7.6); for the latter, the primitive generator is $J_{3}$. The quantum deformation parameter $z$ would introduce an additional perturbation from the non-deformed case $z\to 0$. Expected mathematical and physical properties can be guessed from already known applications of the non-standard deformation, which would be the obtainment of superintegrable Zernike-oscillator chains [54] and the construction of generalized Zernike systems on spaces of non-constant curvature determined by $z$ [55]. We remark that this is still an open issue for the classical Zernike systems.
Work along these various lines is currently in progress.
Acknowledgements
R.C.-S., F.J.H. and A.B. have been partially supported by Agencia Estatal de Investigación (Spain) under the grant PID2023-148373NB-I00 funded by MCIN/AEI/10.13039/501100011033/FEDER, UE. F.J.H. and A.B. also acknowledge support by the Q-CAYLE Project funded by the Regional Government of Castilla y León (Junta de Castilla y León, Spain) and by the Spanish Ministry of Science and Innovation (MCIN) through the European Union funds NextGenerationEU (PRTR C17.I1). The research of D.L. has been partially funded by MUR - Dipartimento di Eccellenza 2023-2027, codice CUP G43C22004580005 - codice progetto DECC23__\__012__\__DIP and partially supported by INFN-CSN4 (Commissione Scientifica Nazionale 4 - Fisica Teorica), MMNLP project. D.L. is a member of GNFM, INdAM. I.M. has been supported by Australian Research Council Future Fellowship FT180100099.